### v1

# On an unconditional $\rm GL_{3}$ analog of Selberg’s result

Abstract.
Let $S_{F}(t)=\pi^{-1}\arg L(1/2+it,F)$, where $F$ is a Hecke–Maass cusp form for $\rm SL_{3}(\mathbb{Z})$ in the generic position with the spectral parameter $\nu_{F}=\big{(}\nu_{F,1},\nu_{F,2},\nu_{F,3}\big{)}$ and the Langlands parameter $\mu_{F}=\big{(}\mu_{F,1},\mu_{F,2},\mu_{F,3}\big{)}$. In this paper, we establish an unconditional asymptotic formula for the moments of $S_{F}(t)$. Previouly, such a formula was only known under the Generalized Riemann Hypothesis. The key ingredient is a weighted zero-density estimate in the spectral aspect for $L(s,F)$ which was recently proved by the authors in [18].
Key words and phrases:
Moments, $\rm GL_{3}$ Hecke–Maass cusp forms, Selberg’s limit theorem.
2010 Mathematics Subject Classification:
11F12, 11F66, 11F72.
Q. Sun was partially supported by the National Natural Science Foundation of China (Grant Nos. 12471005 and 12031008) and the Natural Science Foundation of Shandong Province (Grant No. ZR2023MA003).
1. Introduction
It is a well-established fact that the function $S(t)=\frac{1}{\pi}\arg\zeta\left(\frac{1}{2}+it\right)$ is associated with the number of non-trivial zeros of the Riemann zeta function $\zeta(s)$ up to the height $t>0$. The argument is defined by continuous variation along the horizontal segment from $\infty+it$ to $1/2+it$, commencing at $\infty+it$ with the value 00. The intricate behavior of $S(t)$ represents a highly intriguing topic and has been explored by numerous number theorists (see, e.g., [1, 6, 12, 22]).
In 1940s, Selberg [19, 20] examined the distribution of $S(t)$ and demonstrated that for any $n\in\mathbb{N}$,
$\frac{1}{T}\int_{T}^{2T}S(t)^{2n}\mathrm{d}t=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log T)^{n}+O\big{(}(\log\log T)^{n-1/2}\big{)},$ (1.1)
which implies that, in an average sense, $|S(t)|$ has an order of magnitude $\sqrt{\log\log T}$. Additionally, for the Dirichlet $L$-function $L(s,\chi)$ with primitive $\chi$, Selberg [21] studied the behavior of $S(t,\chi)=\frac{1}{\pi}\arg L\big{(}\frac{1}{2}+it,\chi\big{)}$ and obtained, for prime modulus $q$ and any fixed $t>0$,
$\frac{1}{q}\,\,\sideset{}{{}^{\star}}{\sum}_{\chi\bmod q}S(t,\chi)^{2n}=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log q)^{n}+O\big{(}(\log\log q)^{n-1/2}\big{)},$ (1.2)
where $\star$ indicates that the summation runs over the primitive characters $\chi(\bmod\,q)$. It should be noted that the counterparts of (1.1) and (1.2) for the odd-order moments can also be derived by making minor adjustments to Selberg’s approach.
For the $\rm GL_{2}$ family of $L$-functions, Hejhal–Luo [8] first considered the generalization of (1.2) to the case of eigenfunctions $\psi$ of the hyperbolic Laplacian $\Delta$ for the full modular group. More precisely, let $L(s,\psi)$ be the standard Hecke–Maass $L$-function associated to $\psi$ and define
$S(t,\psi)=\frac{1}{\pi}\arg L\big{(}\frac{1}{2}+it,\psi\big{)}.$
They established the analogue of (1.1) when $\psi$ is an Eisenstain series. And when $\psi$ is the Hecke–Maass cusp form $u_{j}$ of $\rm SL_{2}(\mathbb{Z})$ with the first Fourier coefficient $\nu_{j}(1)$ and Laplacian eigenvalue $\lambda_{j}=\frac{1}{4}+t_{j}^{2}$ ($t_{j}\geq 0$), under Generalized Riemann Hypothesis (GRH), they showed that for fixed $t,\beta>0$ and any large parameter $T>0$,
$\begin{split}\sum_{j=1}^{\infty}e^{-\beta(t_{j}-T)^{2}}|\nu_{j}(1)|^{2}S(t,u_{j})^{2n}=\frac{2\pi^{-3/2}}{\sqrt{\beta}}\frac{(2n)!}{n!(2\pi)^{2n}}T(\log\log T)^{n}+O_{t,n,\beta}\big{(}T(\log\log T)^{n-1/2}\big{)}.\end{split}$
Recently, Sun–Wang [17] reestablished this asymptotic formula by means of the zero-density estimate of weighted type instead of assuming GRH. Moreover, for the holomorphic case, Liu–Shim [14] and Sun–Wang [16] successively obtained unconditional analogues in the weight aspect and in the level aspect, respectively.
For the $\rm GL_{3}$ family of $L$-functions, Liu–Liu [13] explored the average behavior of the analogue function
$S_{F}(t):=\frac{1}{\pi}\arg L(1/2+it,F)$
defined for Hecke–Maass cusp forms $F$ for $\rm SL_{3}(\mathbb{Z})$. To state the result, we need to introduce some notations. Let $F$ run over an orthogonal basis of Hecke–Maass cusp forms for $\rm SL_{3}(\mathbb{Z})$. For each $F$, we let $\mu_{F}=\big{(}\mu_{F,1},\mu_{F,2},\mu_{F,3}\big{)}$ be its Langlands parameter and have that $\mu_{F}$ is a point in the region
$\begin{split}\Lambda_{1/2}^{\prime}=\Bigg{\{}(\mu_{F,1},\mu_{F,2},\mu_{F,3})\in\mathbb{C}^{3},\,\,\,\begin{aligned} &\mu_{F,1}+\mu_{F,2}+\mu_{F,3}=0,\,\,|\operatorname{Re}(\mu_{F,j})|\leq\frac{1}{2},\,j=1,2,3,\\ &\{-\mu_{F,1},-\mu_{F,2},-\mu_{F,3}\}=\{\overline{\mu_{F,1}},\overline{\mu_{F,2}},\overline{\mu_{F,3}}\}\end{aligned}\Bigg{\}}\end{split}$
in the Lie algebra $\mathfrak{sl}_{3}(\mathbb{C})$. Also, the cusp form $F$ has the spectral parameter $\nu_{F}=\big{(}\nu_{F,1},\nu_{F,2},\nu_{F,3}\big{)}$, where
$\begin{split}\nu_{F,1}:=\frac{1}{3}&\big{(}\mu_{F,1}-\mu_{F,2}\big{)},\quad\nu_{F,2}:=\frac{1}{3}\big{(}\mu_{F,2}-\mu_{F,3}\big{)},\\ \nu_{F,3}:=&-\nu_{F,1}-\nu_{F,2}=\frac{1}{3}\big{(}\mu_{F,3}-\mu_{F,1}\big{)}.\end{split}$ (1.3)
And we have
$\mu_{F,1}=2\nu_{F,1}+\nu_{F,2},\quad\mu_{F,2}=\nu_{F,2}-\nu_{F,1},\quad\mu_{F,3}=-\nu_{F,1}-2\nu_{F,2}.$ (1.4)
We will use the coordinates $\mu$ and $\nu$ simultaneously. Note that $\big{(}\nu_{F,1}-\frac{1}{3},\nu_{F,2}-\frac{1}{3}\big{)}$ is the type of $F$ as defined in Goldfeld’s book [7]. Assume that $F$ has the normalized Fourier coefficients $A_{F}(m,n)$ with $A_{F}(1,1)=1$. The $L$-function associated with $F$ is defined by
$L(s,F)=\sum_{n=1}^{\infty}\frac{A_{F}(1,n)}{n^{s}},\quad\,\operatorname{Re}(s)>1,$
and has the Euler product
$\begin{split}L(s,F)&=\prod_{p}\left(1-\frac{A_{F}(1,p)}{p^{s}}+\frac{A_{F}(p,1)}{p^{2s}}-\frac{1}{p^{3s}}\right)^{-1}\\ &=\prod_{p}\big{(}1-\alpha_{F,1}(p)p^{-s}\big{)}^{-1}\big{(}1-\alpha_{F,2}(p)p^{-s}\big{)}^{-1}\big{(}1-\alpha_{F,3}(p)p^{-s}\big{)}^{-1}.\end{split}$
Here we have
$\begin{split}\begin{cases}\alpha_{F,1}(p)\alpha_{F,2}(p)\alpha_{F,3}(p)=1\\ \alpha_{F,1}(p)+\alpha_{F,2}(p)+\alpha_{F,3}(p)=A_{F}(1,p)\\ \alpha_{F,1}(p)\alpha_{F,2}(p)+\alpha_{F,2}(p)\alpha_{F,3}(p)+\alpha_{F,3}(p)\alpha_{F,1}(p)=A_{F}(p,1),\end{cases}\end{split}$ (1.5)
and by Kim–Sarnak [10] that
$\begin{split}|\alpha_{F,j}(p)|\leq p^{\frac{5}{14}},\quad\text{and}\quad|\operatorname{Re}(\mu_{F,j})|\leq\frac{5}{14},\quad\text{for}\,\,j=1,2,3.\end{split}$ (1.6)
Note that $L(s,F)$ has an analytic continuation to the whole complex plane and satisfies the functional equation
$\Lambda(s,F)=L(s,F)\prod_{j=1}^{3}\Gamma_{\mathbb{R}}(s-\mu_{F,j})=L(1-s,\widetilde{F})\prod_{j=1}^{3}\Gamma_{\mathbb{R}}(1-s+\mu_{F,j})=\Lambda(1-s,\widetilde{F}),$
where $\Gamma_{\mathbb{R}}(s)=\pi^{-s/2}\Gamma(s/2)$ and the dual form of $F$ is denoted by $\widetilde{F}$ with the Langlands parameter $(-\mu_{F,1},-\mu_{F,2},-\mu_{F,3})$, the spectral parameter $\big{(}\nu_{F,2},\nu_{F,1},\nu_{F,3}\big{)}$ and the coefficients $A_{\widetilde{F}}(m,n)=\overline{A_{F}(m,n)}=A_{F}(n,m)$.
Let $\mu_{0}=(\mu_{0,1},\mu_{0,2},\mu_{0,3})$ and $\nu_{0}=(\nu_{0,1},\nu_{0,2},\nu_{0,3})$, and satisfy the corresponding relations (1.3) and (1.4). We also assume
$|\mu_{0,j}|\asymp|\nu_{0,j}|\asymp T:=\|\mu_{0}\|,\quad 1\leq j\leq 3,$ (1.7)
where
$\|\mu_{0}\|=\sqrt{|\mu_{0,1}|^{2}+|\mu_{0,2}|^{2}+|\mu_{0,3}|^{2}}.$
Throughout this paper, we let $M\!=\!T^{1-\eta}$ for some very small $20<\eta<1/$ 2. Define a test function $h_{T,M}(\mu)$ (depending on $\mu_{0}$) for $\mu=(\mu_{1},\mu_{2},\mu_{3})\in\mathbb{C}^{3}$ by
$h_{T,M}(\mu):=P(\mu)^{2}\bigg{(}\sum_{w\in\mathcal{W}}\psi\bigg{(}\frac{w(\mu)-\mu_{0}}{M}\bigg{)}\bigg{)}^{2},$
where $\psi(\mu)=\exp\left(\mu_{1}^{2}+\mu_{2}^{2}+\mu_{3}^{2}\right)$ and
$P(\mu)=\prod_{0\leq n\leq A}\prod_{j=1}^{3}\frac{\left(\nu_{j}-\frac{1}{3}(1+2n)\right)\left(\nu_{j}+\frac{1}{3}(1+2n)\right)}{|\nu_{0,j}|^{2}}$
for some fixed large $A>0$. Here
$\mathcal{W}:=\left\{I,\;w_{2}=\left(\begin{smallmatrix}1&&\\ &&-1\\ &1&\end{smallmatrix}\right),\;w_{3}=\left(\begin{smallmatrix}&1&\\ -1&&\\ &&1\end{smallmatrix}\right),\;w_{4}=\left(\begin{smallmatrix}&1&\\ &&1\\ 1&&\end{smallmatrix}\right),\;w_{5}=\left(\begin{smallmatrix}&&1\\ 1&&\\ &1&\end{smallmatrix}\right),\;w_{6}=\left(\begin{smallmatrix}&&1\\ &-1&\\ 1&&\end{smallmatrix}\right)\right\}$
is the Weyl group of $\rm SL_{3}(\mathbb{R})$. The function $h_{T,M}(\mu)$ localizes at a ball of radius $M$ about $w(\mu_{0})$ for each $w\in\mathcal{W}$. Moreover, we define the spectral measure on the hyperplane $\mu_{1}+\mu_{2}+\mu_{3}=0$ by
$\mathrm{d}_{\text{spec}}\mu=\text{spec}(\mu)\mathrm{d}\mu,$
where
$\text{spec}(\mu):=\prod_{j=1}^{3}\left(3\nu_{j}\tan\left(\frac{3\pi}{2}\nu_{j}\right)\right)\quad\mbox{and}\quad\mathrm{d}\mu=\mathrm{d}\mu_{1}\mathrm{d}\mu_{2}=\mathrm{d}\mu_{2}\mathrm{d}\mu_{3}=\mathrm{d}\mu_{1}\mathrm{d}\mu_{3}.$
By trivial estimate, we have
$\mathcal{H}:=\frac{1}{192\pi^{5}}\int_{\operatorname{Re}(\mu)=0}h_{T,M}(\mu)\mathrm{d}_{\rm{spec}}\mu\asymp T^{3}M^{2}.$ (1.8)
Fix any $t>0$ and $n\in\mathbb{N}$, under GRH, Liu and Liu [13] established that, for large positive $T$,
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\nu_{F})}{L(1,\text{Ad}^{2}F)}S_{F}(t)^{2n}=\frac{9(2n)!}{4n!(2\pi)^{2n}}(\log\log T)^{n}+O_{t,n}\big{(}(\log\log T)^{n-1/2}\big{)},\end{split}$ (1.9)
where $L(1,\text{Ad}^{2}F)$ denotes the adjoint square $L$-function of $F$. In this paper, we are devoted to proving the asymptotic formula (1.9) without assuming the GRH. We introduce the normalizing factor
$\mathcal{N}_{F}:=\|F\|^{2}\prod_{j=1}^{3}\cos\left(\frac{3}{2}\pi\nu_{F,j}\right).$
By Rankin–Selberg theory in combination with Stade’s formula (see e.g., Blomer [3, Section 4] and Li [11, Theorem 2]), it is easy to see that
$\begin{split}\mathcal{N}_{F}\asymp\mathop{\operatorname{Res}}_{s=1}L(s,F\times\widetilde{F})\ll\|\mu_{F}\|^{\varepsilon},\end{split}$
where $L(s,F\times\widetilde{F})$ is the Rankin–Selberg square of $F$ and the implied constants depending at most on $\varepsilon$. Our main result is the following theorem.
Theorem 1.1.
Let $t>0$ be fixed and $T$ be any large parameter, we have
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}S_{F}(t)^{n}=C_{n}(\log\log T)^{n/2}+O_{t,n}\big{(}(\log\log T)^{(n-1)/2}\big{)},\end{split}$ (1.10)
where
$C_{n}=\begin{cases}\frac{n!}{(n/2)!(2\pi)^{n}},\,&\textit{if }\,n\,\,\text{is even},\\ 0,\,&\textit{if }\,n\,\,\text{is odd}.\end{cases}$
Moreover, we define the following probability measure $\mu_{TM}$ on $\mathbb{R}$ by
$\begin{split}\mu_{TM}(E):=\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\mathbf{1}_{E}\left(\frac{S_{F}(t)}{\sqrt{\log\log T}}\right)\right)\bigg{/}\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\right),\end{split}$
where $\mathbf{1}_{E}$ is the characteristic function of a Borel measurable set $E$ in $\mathbb{R}$. Then for every $a<b$, we have
$\lim_{T\rightarrow\infty}\mu_{TM}([a,b])=\sqrt{\pi}\int_{a}^{b}\exp(-\pi^{2}\xi^{2})\mathrm{d}\xi,$
uniformly with respect to $M$.
Remark 1.
For $F$ a Hecke–Maass cusp form of $\rm SL_{3}(\mathbb{Z})$, on an average sense, our result indicates that $|S_{F}(t)|$ has order of magnitude $\sqrt{\log\log T}$ under the assumption (1.7). In addition, as $T\rightarrow\infty$, the ratios $S_{F}(t)/\sqrt{\log\log T}$ converges in distribution to a normal distribution of mean 00 and variance $(2\pi^{2})^{-1}$.
Notation Throughout the paper, $\varepsilon$ is an arbitrarily small positive real number, which may be different at each occurrence. As usual, $e(z)=\exp(2\pi iz)=e^{2\pi iz}$ and $\delta_{m,n}$ is the Kronecker symbol that detects $m=n$. We write $f=O(g)$ or $f\ll g$ to mean $|f|\leq cg$ for some unspecified positive constant $c$. The symbol $y\asymp Y$ is to mean that $c_{1}Y\leq|y|\leq c_{2}Y$ for some positive constants $c_{1}$ and $c_{2}$.
Structure of this paper. This paper is organized as follows. In Sect. 2, we give some preliminary results needed for the proof. In Sect. 3, we establish some lemmas and provide a crucial approximation of $S_{F}(t)$. In Sect. 4, we give the detailed proof of Theorem 1.1.
2. Preliminaries
2.1. Hecke relations
Firstly, let’s recall the Hecke relations (see Goldfeld [7, pg. 168]):
$\begin{split}A_{F}(m_{1}m_{2},n_{1}n_{2})&=A_{F}(m_{1},n_{1})A_{F}(m_{2},n_{2}),\quad\text{if}\,\,(m_{1}n_{1},m_{2}n_{2})=1,\\ A_{F}(n,1)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\left(\frac{m_{1}d_{0}}{d_{1}},\frac{m_{2}d_{1}}{d_{2}}\right),\\ A_{F}(1,n)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\left(\frac{m_{1}d_{2}}{d_{1}},\frac{m_{2}d_{0}}{d_{2}}\right),\\ A_{F}(n_{1},1)A_{F}(1,n_{2})&=\sum_{d|(n_{1},n_{2})}A_{F}\left(\frac{n_{1}}{d},\frac{n_{2}}{d}\right).\end{split}$
Then we arrive at the following useful result, which was presented by Liu–Liu in [13, Proposition 5.1].
Lemma 2.1.
For each prime $p$ and each integer $n\geq 0$, we have
$\begin{split}A_{F}(1,p)^{n}=\sum_{k,\ell\geq 0\atop k+\ell\leq n}a_{k,\ell}(n)A_{F}(p^{k},p^{\ell})\end{split}$
for some integers $a_{k,\ell}(n)\geq 0$. Furthermore, $a_{k,\ell}(n)\neq 0$ if and only if $(k,\ell)$ belongs to
$\begin{split}E_{n}:=\big{\{}(n_{2}-n_{1},n_{3}-n_{2})\,|\,n_{j}\in\mathbb{Z},n_{3}\geq n_{2}\geq n_{1}\geq 0,n_{1}+n_{2}+n_{3}=n\big{\}}.\end{split}$
2.2. A summation formula
We shall need the following summation formula, which is an application of the $\rm GL_{3}$ Kuznetsov trace formula in the version of Buttcane [4, Theorems 2–4].
Theorem 2.2.
Let $P=m_{1}n_{1}m_{2}n_{2}\neq 0$. We have for any $\varepsilon>0$ that
$\begin{split}&\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\overline{A_{F}(m_{1},m_{2})}A_{F}(n_{1},n_{2})\\ &\quad=\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}\mathcal{H}+O\left((TP)^{\varepsilon}\big{(}TP^{1/2}+T^{3}+TM^{2}P^{\vartheta}\big{)}\right),\end{split}$
where $\vartheta=7/64$ is the best result towards the Ramanujan–Petersson conjecture for Fourier coefficients of $\rm GL_{2}$ Maass cusp forms due to Kim–Sarnak [10].
Proof.
See Sun–Wang [18, Theorem 7.1]. ∎
We record here the frequently used estimate
$\mathcal{H}=CT^{3}M^{2}+O(T^{2}M^{3}),$
for some constant $C>0$ (see Buttcane–Zhou [5, pg. 15]). Taking $m_{1}=n_{1}=m_{2}=n_{2}=1$ in Theorem 2.2 immediately gives
Corollary 2.3.
With the assumptions in Theorem 2.2. We have
$\begin{split}&\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}=CT^{3}M^{2}+O(T^{2}M^{3}).\end{split}$
2.3. The zero-density estimate
We will need the following weighted version of zero-density estimate in the spectral aspect for the $L$-functions associated to $\rm GL_{3}$ Hecke–Maass cusp forms which was established by Sun–Wang [18, Theorem 1.1].
Lemma 2.4.
For $\sigma>0$ and $H>0$, we let
$\begin{split}N(\sigma,H;F):=\#\big{\{}\rho=1/2+\beta+i\gamma:L(\rho,F)=0,\beta>\sigma,|\gamma|<H\big{\}}\end{split}$
and
$\begin{split}N(\sigma,H):=\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}N(\sigma,H;F).\end{split}$
Let $22/\log T<\sigma<1/$ 2. For some sufficiently small $\delta_{1},\,\theta>0$, we have
$\begin{split}N(\sigma,H)\ll HT^{-\theta\sigma}\log T,\end{split}$
uniformly in $3/\log T<H<T^{\delta_{1}}$.
3. An approximation of $S_{F}(t)$
In this section we will prove several lemmas and derive an approximation of $S_{F}(t)$. We denote by $\rho=\beta+i\gamma$ a typical zero of $L(s,F)$ inside the critical strip, i.e., $10<\beta<$ 1. For $\operatorname{Re}(s)>1$, we have the Dirichlet series expansion for the logarithmic derivative of $L(s,F)$,
$-\frac{L^{\prime}}{L}(s,F)=\sum_{n=1}^{\infty}\frac{\Lambda(n)C_{F}(n)}{n^{s}},$ (3.1)
where $\Lambda(n)$ denotes the von Mangoldt function, and
$C_{F}(n)=\begin{cases}\alpha_{F,1}(p)^{k}+\alpha_{F,2}(p)^{k}+\alpha_{F,3}(p)^{k},\,&\text{if}\,\,n=p^{k}\,\,\text{for a prime}\,\,p,\\ 0,\,&\text{otherwise}.\end{cases}$ (3.2)
Lemma 3.1.
Let $x>1$. For $s\neq\rho$, and $s\neq-2m+\mu_{F,j}$, ($m\geq 0$, $j=1,2,3$), we have the following identity
$\begin{split}\frac{L^{\prime}}{L}(s,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{s}}-\frac{1}{\log^{2}x}\sum_{\rho}\frac{x^{\rho-s}(1-x^{\rho-s})^{2}}{(\rho-s)^{3}}\\ &-\frac{1}{\log^{2}x}\sum_{j=1}^{3}\sum_{m=0}^{\infty}\frac{x^{-2m-s+\mu_{F,j}}(1-x^{-2m-s+\mu_{F,j}})^{2}}{(-2m-s+\mu_{F,j})^{3}},\end{split}$
where
$\Lambda_{x}(n)=\begin{cases}\Lambda(n),\,&\textit{if }\,\,n\leq x,\\ \Lambda(n)\frac{\log^{2}(x^{3}/n)-2\log^{2}(x^{2}/n)}{2\log^{2}x},\,&\textit{if }\,\,x\leq n\leq x^{2},\\ \Lambda(n)\frac{\log^{2}(x^{3}/n)}{2\log^{2}x},\,&\textit{if }\,\,x^{2}\leq n\leq x^{3},\\ 0,\,&\textit{if }\,\,n\geq x^{3}.\\ \end{cases}$
Proof.
Recall the discontinuous integral
$\frac{1}{2\pi i}\int_{(\alpha)}\frac{y^{s}}{s^{3}}\mathrm{d}s=\begin{cases}\frac{\log^{2}y}{2},\,&\textit{if }\,\,y\geq 1,\\ 0,\,&\textit{if }\,\,0<y\leq 1\end{cases}$
for $\alpha>0$. It follows from (3.1) and (3.2) that
$-\log^{2}x\sum_{n=1}^{\infty}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{s}}=\frac{1}{2\pi i}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v,$
where $\alpha=\max\{2,1+\operatorname{Re}(s)\}$. By moving the line of integration all way to the left, we pick up the residues at $v=0$, $v=\rho-s$ and $v=-2m-s+\mu_{F,j}$ ($m\geq 0$, $j=1,2,3$), and deduce that
$\begin{split}&\frac{1}{2\pi i}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v\\ =&\frac{L^{\prime}}{L}(s,F)\log^{2}x+\sum_{\rho}\frac{x^{\rho-s}(1-x^{\rho-s})^{2}}{(\rho-s)^{3}}+\sum_{j=1}^{3}\sum_{m=0}^{\infty}\frac{x^{-2m-s+\mu_{F,j}}(1-x^{-2m-s+\mu_{F,j}})^{2}}{(-2m-s+\mu_{F,j})^{3}}.\end{split}$
Thus we complete the proof of lemma. ∎
Lemma 3.2.
For $s=\sigma+it$, $s^{\prime}=\sigma^{\prime}+it^{\prime}$ such that $1/2\leq\sigma,\sigma^{\prime}\leq 2$, $s\neq\rho$ and $s^{\prime}\neq\rho$, we have
$\operatorname{Im}\left(\frac{L^{\prime}}{L}(s,F)-\frac{L^{\prime}}{L}(s^{\prime},F)\right)=\operatorname{Im}\sum_{\rho}\left(\frac{1}{s-\rho}-\frac{1}{s^{\prime}-\rho}\right)+O(1),$
and
$\operatorname{Re}\frac{L^{\prime}}{L}(s,F)=\sum_{\rho}\frac{\sigma-\beta}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}+O(\log(|t|+M_{F})),$ (3.3)
where
$\begin{split}M_{F}:=2+\max_{1\leq j\leq 3}|\mu_{F,j}|.\end{split}$
Proof.
By the Hadamard factorization theorem of the entire function $\Lambda(s,F)$, we have
$\frac{L^{\prime}}{L}(s,F)=b_{F}+\sum_{\rho}\left(\frac{1}{s-\rho}+\frac{1}{\rho}\right)-\frac{1}{2}\sum_{j=1}^{3}\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{s-\mu_{F,j}}{2}\right)+\frac{3}{2}\log\pi,$
for some $b_{F}\in\mathbb{C}$ with $\operatorname{Re}(b_{F})=-\operatorname{Re}\sum_{\rho}\frac{1}{\rho}$ (see [9, Proposition 5.7 (3)]). Note that for $z\not\in-\mathbb{N}$,
$\frac{\Gamma^{\prime}}{\Gamma}(z)=-\gamma+\sum_{k\geq 1}\left(\frac{1}{k}-\frac{1}{k-1+z}\right),$ (3.4)
where $\gamma$ is the Euler constant. Using (3.4), we get that for $\frac{1}{4}\leq u$,
$\operatorname{Im}\frac{\Gamma^{\prime}}{\Gamma}(u+iv)=\sum_{k\geq 1}\frac{v}{(k+u-1)^{2}+v^{2}}\ll 1.$
Combined with the asymptotic formula of the logarithmic derivative of $\Gamma(z)$,
$\frac{\Gamma^{\prime}}{\Gamma}(z)=\log z-\frac{1}{2z}+O_{\varepsilon}\left(\frac{1}{|z|^{2}}\right),\quad\text{for}\,\,|\arg z|\leq\pi-\varepsilon,$
we complete the proof of lemma. ∎
Let $x\geq 4$. We define
$\sigma_{x}=\sigma_{x,F}=\sigma_{x,F,t}:=\frac{1}{2}+2\max_{\rho}\left\{\left|\beta-\frac{1}{2}\right|,\frac{5}{\log x}\right\},$ (3.5)
where $\rho=\beta+i\gamma$ runs through the zeros of $L(s,F)$ for which
$|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}.$
The following lemma is to display the dependence of $\sigma_{x}$ as needed.
Lemma 3.3.
Let $x\geq 4$. For $\sigma\geq\sigma_{x}$, we have
$\begin{split}\frac{L^{\prime}}{L}(\sigma+it,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}+O\left(x^{1/4-\sigma/2}\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)\\ &+O\left(x^{1/4-\sigma/2}\log(|t|+M_{F})\right),\end{split}$
and
$\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}=O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\big{(}\log(|t|+M_{F})\big{)}.$
Proof.
By (3.3), we have
$\operatorname{Re}\frac{L^{\prime}}{L}(\sigma_{x}+it,F)=\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\log(|t|+M_{F})).$ (3.6)
Moreover, if $\rho=\beta+i\gamma$ is zero of $L(s,F)$, then $(1-\beta)+i\gamma$ is also a zero of $L(s,F)$. Thus we have
$\begin{split}&\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\\ =&\frac{1}{2}\left(\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+\sum_{\rho}\frac{\sigma_{x}-(1-\beta)}{(\sigma_{x}-(1-\beta))^{2}+(t-\gamma)^{2}}\right)\\ =&\left(\sigma_{x}-\frac{1}{2}\right)\sum_{\rho}\frac{(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})((\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2})}.\end{split}$ (3.7)
If $|\beta-1/2|\leq\frac{\sigma_{x}-1/2}{2}$, then
$\begin{split}(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}&\geq\frac{1}{2}\big{(}(\sigma_{x}-1/2)^{2}+(\beta-1/2)^{2}\big{)}\\ &=\frac{1}{4}\big{(}(\sigma_{x}-\beta)^{2}+(\sigma_{x}-1+\beta)^{2}\big{)}.\end{split}$
Thus,
$(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}\geq\frac{1}{4}\big{(}(\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2}\big{)}.$ (3.8)
If $|\beta-1/2|>\frac{\sigma_{x}-1/2}{2}$, then by (3.5) and (3) we have
$|t-\gamma|>\frac{x^{3|\beta-1/2|}}{\log x}>3|\beta-1/2|.$
Thus,
$\begin{split}&(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}\\ &=\big{(}(\sigma_{x}-1/2)^{2}+(\beta-1/2)^{2}\big{)}+(t-\gamma)^{2}-2(\beta-1/2)^{2}\\ &\geq\frac{1}{2}\big{(}(\sigma_{x}-\beta)^{2}+(\sigma_{x}-1+\beta)^{2}\big{)}+\frac{7}{9}(t-\gamma)^{2}\\ &\geq\frac{1}{4}\big{(}(\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2}\big{)}.\end{split}$ (3.9)
Combining (3.8), (3.9) and (3.7), we have
$\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\geq\frac{1}{4}(\sigma_{x}-1/2)\sum_{\rho}\frac{1}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})}.$
Using this bound and (3.6), we obtain
$\sum_{\rho}\frac{1}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})}\leq\frac{4}{\sigma_{x}-1/2}\left|\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\right|+O\left(\frac{\log(|t|+M_{F})}{\sigma_{x}-1/2}\right).$ (3.10)
On the other hand, by Lemma 3.1 and the bound (1.6), we have
$\begin{split}\frac{L^{\prime}}{L}(\sigma+it,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}+\frac{w_{F}(x,\sigma,t)}{\log^{2}x}\sum_{\rho}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big{(}(\sigma-\beta)^{2}+(t-\gamma)^{2}\big{)}^{3/2}}+O\left(\frac{x^{5/14-\sigma}}{\log^{2}x}\right),\end{split}$ (3.11)
with $|w_{F}(x,\sigma,t)|\leq 1$. Next we claim that
$\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big{(}(\sigma-\beta)^{2}+(t-\gamma)^{2}\big{)}^{3/2}}\leq 2\log x\frac{x^{1/4-\sigma/2}}{(\beta-\sigma_{x})^{2}+(t-\gamma)^{2}}.$ (3.12)
If $\beta\leq\frac{\sigma_{x}+1/2}{2}$, then we have
$\begin{split}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big{(}(\sigma-\beta)^{2}+(t-\gamma)^{2}\big{)}^{3/2}}&\leq\frac{4x^{1/4-\sigma/2}}{(\sigma_{x}-\beta)\big{(}(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big{)}}\\ &\leq\frac{8}{\sigma_{x}-1/2}\frac{x^{1/4-\sigma/2}}{\big{(}(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big{)}}\\ &\leq\frac{4}{5}\log x\frac{x^{1/4-\sigma/2}}{\big{(}(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big{)}}.\end{split}$
If $\beta>\frac{\sigma_{x}+1/2}{2}$, then we have
$|t-\gamma|>\frac{x^{3|\beta-1/2|}}{\log x}>3|\beta-1/2|\geq 3|\beta-\sigma_{x}|.$
Thus, $(t-\gamma)^{2}>\frac{8}{9}\big{(}(\beta-\sigma_{x})^{2}+(t-\gamma)^{2}\big{)}$. Hence
$\begin{split}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big{(}(\sigma-\beta)^{2}+(t-\gamma)^{2}\big{)}^{3/2}}&\leq\frac{x^{\beta-\sigma}(1+x^{\beta-1/2})^{2}}{|t-\gamma|(t-\gamma)^{2}}\leq\frac{\log x}{x^{3|\beta-1/2|}}\frac{9}{8}\frac{x^{\beta-\sigma}(1+x^{\beta-1/2})^{2}}{(\beta-\sigma_{x})^{2}+(t-\gamma)^{2}}\\ &=\frac{9}{8}(\log x)(1+x^{-(\beta-1/2)})^{2}\frac{x^{1/2-\sigma}}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\\ &\leq\frac{9}{8}(\log x)(1+e^{-5})^{2}\frac{x^{1/2-\sigma}}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\\ &<2(\log x)\frac{x^{1/2-\sigma}}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}.\end{split}$
Using (3.10) and (3.12), we have
[…]
Inserting this bound into (3.11), we have
$\begin{split}\frac{L^{\prime}}{L}(\sigma+it,F)\leq&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}+\frac{4}{5}w^{\prime}_{F}(x,\sigma,t)x^{1/4-\sigma/2}\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\\ &+O\left(x^{1/4-\sigma/2}\log(|t|+M_{F})\right),\end{split}$ (3.13)
with $|w^{\prime}_{F}(x,\sigma,t)|\leq 1$. By taking $\sigma=\sigma_{x}$,
$\begin{split}&\left(1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\right)\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\\ =&\,O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\left(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\right).\end{split}$
Since $\left|1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\right|\geq 1-\left|\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\right|\geq\frac{1}{5}$, we have
$\frac{L^{\prime}}{L}(\sigma_{x}+it,F)=O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\left(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\right).$ (3.14)
Inserting (3.14) into (3.13) and (3.10), respectively, we complete the proof of lemma. ∎
The following theorem is to provide an approximation of $S_{F}(t)$.
Theorem 3.4.
For $t\neq 0$, and $x\geq 4$, we have
$\begin{split}S_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}\log n}+O\left((\sigma_{x}-1/2)\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)\\ &+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)},\end{split}$
where $\sigma_{x}$ is defined in (3.5).
Proof.
By the definition of $S_{F}(t)$, we have
$\begin{split}\pi S_{F}(t)=&-\int_{1/2}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+it,F)\mathrm{d}\sigma\\ =&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+it,F)\mathrm{d}\sigma-(\sigma_{x}-1/2)\operatorname{Im}\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\\ &+\int_{1/2}^{\sigma_{x}}\operatorname{Im}\left(\frac{L^{\prime}}{L}(\sigma_{x}+it,F)-\frac{L^{\prime}}{L}(\sigma+it,F)\right)\mathrm{d}\sigma\\ :=&\,J_{1}+J_{2}+J_{3}.\end{split}$
By Lemma 3.3, we have
$\begin{split}J_{1}=&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+it,F)\mathrm{d}\sigma\\ =&\int_{\sigma_{x}}^{\infty}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}\mathrm{d}\sigma+O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\right)\\ &+O\left(\log(|t|+M_{F})\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\right)\\ =&\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}\log n}+O\left(\frac{1}{\log x}\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\left(\frac{\log(|t|+M_{F})}{\log x}\right).\end{split}$
Taking $\sigma=\sigma_{x}$ in Lemma 3.3, we have
$\begin{split}J_{2}\leq&\,(\sigma_{x}-1/2)\left|\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\right|\\ \ll&\,(\sigma_{x}-1/2)\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|+(\sigma_{x}-1/2)\log(|t|+M_{F}).\end{split}$
Using Lemma 3.2, we get
$\begin{split}&\operatorname{Im}\left(\frac{L^{\prime}}{L}(\sigma_{x}+it,F)-\frac{L^{\prime}}{L}(\sigma+it,F)\right)\\ =&\sum_{\rho}\frac{(\sigma_{x}-\sigma)(\sigma+\sigma_{x}-2\beta)(t-\gamma)}{\big{(}(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big{)}\big{(}(\sigma-\beta)^{2}+(t-\gamma)^{2}\big{)}}+O(1).\end{split}$
Hence
$\begin{split}|J_{3}|\leq&\sum_{\rho}\int_{1/2}^{\sigma_{x}}\frac{(\sigma_{x}-\sigma)|\sigma+\sigma_{x}-2\beta||t-\gamma|}{\big{(}(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big{)}\big{(}(\sigma-\beta)^{2}+(t-\gamma)^{2}\big{)}}\mathrm{d}\sigma+O(\sigma_{x}-1/2)\\ \leq&\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma+O(\sigma_{x}-1/2).\end{split}$
If $|\beta-1/2|\leq\frac{1}{2}(\sigma_{x}-1/2)$, then for $1/2\leq\sigma\leq\sigma_{x}$,
$\begin{split}|\sigma+\sigma_{x}-2\beta|&=|(\sigma-1/2)+(\sigma_{x}-1/2)-2(\beta-1/2)|\\ &\leq|\sigma-1/2|+|\sigma_{x}-1/2|+2|\beta-1/2|\leq 3(\sigma_{x}-1/2).\end{split}$
Thus,
$\begin{split}\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma&\leq 3(\sigma_{x}-1/2)\int_{-\infty}^{\infty}\frac{|t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma\\ &\leq 10(\sigma_{x}-1/2).\end{split}$ (3.15)
If $|\beta-1/2|>\frac{1}{2}(\sigma_{x}-1/2)$, then
$|t-\gamma|>\frac{x^{3|\beta-1/2|}}{\log x}>3|\beta-1/2|$
and for $1/2\leq\sigma\leq\sigma_{x}$,
$\begin{split}|\sigma+\sigma_{x}-2\beta|\leq|\sigma-1/2|+|\sigma_{x}-1/2|+2|\beta-1/2|\leq 6|\beta-1/2|.\end{split}$
Thus,
$\begin{split}\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma&<\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta|}{|t-\gamma|}\mathrm{d}\sigma\\ &\leq\int_{1/2}^{\sigma_{x}}\frac{6|\beta-1/2|}{3|\beta-1/2|}\mathrm{d}\sigma=2(\sigma_{x}-1/2).\end{split}$ (3.16)
By the bounds (3.15)–(3.16) and Lemma 3.3, we obtain
$\begin{split}|J_{3}|\leq&10(\sigma_{x}-1/2)\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\sigma_{x}-1/2)\\ =&O\left((\sigma_{x}-1/2)\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}.\end{split}$
Thus we complete the proof of lemma. ∎
4. Proof of Theorem 1.1
Lemma 4.1.
Let $t>0$ be fixed and $T$ be any large parameter. For $x=T^{\delta/3}$ with $0<\delta<\frac{3\theta}{8nk+3}$, $k\in\mathbb{N}$, we have
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{4n}x^{4nk(\sigma_{x}-1/2)}\ll_{t,n,\delta,k}\frac{1}{(\log T)^{4n}},\end{split}$
where $\sigma_{x,F}$ is defined in (3.5) and $\theta$ is as in Lemma 2.4.
Proof.
For $k\in\mathbb{N}$, by the definition of $\sigma_{x,F}$, we have
$\begin{split}&(\sigma_{x,F}-1/2)^{4n}x^{4nk(\sigma_{x,F}-1/2)}\\ \leq&\left(\frac{10}{\log x}\right)^{4n}x^{40nk/\log x}+2^{4n+1}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}.\end{split}$
On the other hand,
$\begin{split}&\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\left(\frac{i+1}{\log x}\right)^{4n}x^{8nk\frac{i+1}{\log x}}\sum_{\frac{1}{2}+\frac{i}{\log x}<\beta\leq\frac{1}{2}+\frac{i+1}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}1\\ \leq&\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\left(i+1\right)^{4n}e^{8nk(i+1)}N\left(\frac{i}{\log x},|t|+\frac{e^{3(i+1)}}{\log x};F\right).\end{split}$
By Lemma 2.4, we have
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\left(i+1\right)^{4n}e^{8nk(i+1)}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}N\left(\frac{i}{\log x},|t|+\frac{e^{3(i+1)}}{\log x};F\right)\\ \ll&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\infty}\left(i+1\right)^{4n}e^{8nk(i+1)}\left(|t|+\frac{e^{3(i+1)}}{\log x}\right)T^{-\theta\frac{i}{\log x}}\log T\\ \ll&{}_{t,n,\delta,k}\,\frac{\log T}{(\log x)^{4n+1}}\sum_{i=5}^{\infty}\left(i+1\right)^{4n}e^{(8nk+3-\frac{3\theta}{\delta})i}\\ \ll&{}_{t,n,\delta,k}\,\frac{1}{(\log T)^{4n}}\end{split}$
provided that $0<\delta<\frac{3\theta}{8nk+3}$. In addition, by Corollary 2.3,
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\frac{10}{\log x}\right)^{4n}x^{40nk/\log x}\ll_{n,k}\frac{1}{(\log x)^{4n}}\ll_{n,k,\delta}\frac{1}{(\log T)^{4n}}.\end{split}$
Thus we complete the proof of lemma. ∎
For a positive parameter $x$ (to be chosen later), let
$\mathcal{M}_{F}(t):=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+it}}\qquad\text{and}\qquad\mathcal{R}_{F}(t):=S_{F}(t)-\mathcal{M}_{F}(t),$ (4.1)
where we recall (3.2) that $C_{F}(p)=\alpha_{F,1}(p)+\alpha_{F,2}(p)+\alpha_{F,3}(p)=A_{F}(1,p)$.
Lemma 4.2.
We have
$\begin{split}\mathcal{R}_{F}(t)=&O\left(\left|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}\right|\right)+O\left(\sum_{k=2}^{7}\left|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}\right|\right)\\ &+O\left((\sigma_{x}-1/2)\sum_{k=1}^{7}x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)\\ &+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}+O\big{(}1\big{)}.\end{split}$
Proof.
By Theorem 3.4,
$\begin{split}\mathcal{R}_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)(\Lambda_{x}(p)p^{1/2-\sigma_{x}}-\Lambda(p))}{p^{1/2+it}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}\log p^{k}}\\ &+O\left((\sigma_{x}-1/2)\left|\sum_{k=1}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}\right|\right)+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}.\end{split}$
Using the bound (1.6), we deduce that
$\begin{split}\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}=O(1),\quad\text{and}\quad\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}\log p}=O(1).\end{split}$
Thus,
$\begin{split}\mathcal{R}_{F}(t)=&\,\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p^{k}}\\ &+\frac{1}{\pi}\operatorname{Im}\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+it)}\log p^{k}}\\ &+O\left((\sigma_{x}-1/2)\left|\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}\right|\right)\\ &+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}+O(1).\end{split}$
Note that for $1/2\leq a\leq\sigma_{x}$ and $k\geq 1$,
$\begin{split}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+it)}}\right|&=kx^{k(a-1/2)}\left|\int_{a}^{\infty}x^{k(1/2-\sigma)}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\mathrm{d}\sigma\right|\\ &\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma.\end{split}$
Thus,
$\begin{split}&\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+it)}\log p^{k}}\right|=\left|\int_{1/2}^{\sigma_{x}}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+it)}}\mathrm{d}a\right|\\ &\leq k(\sigma_{x}-1/2)x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma,\end{split}$
and
$\begin{split}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}\right|\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma.\end{split}$
Thus we complete the proof of lemma. ∎
Lemma 4.3.
Let $t>0$ be fixed and $T$ be any large parameter. For $x=T^{\delta/3}$ with sufficiently small $\delta>0$, we have
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\mathcal{M}_{F}(t)^{n}=C_{n}(\log\log T)^{\frac{n}{2}}+O_{t,n}\big{(}(\log\log T)^{\frac{n-1}{2}}\big{)}\end{split}$ (4.2)
and
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}=O_{t,n}(1),\end{split}$ (4.3)
where the $O$-term depends only on $t$ and $n$, $C_{n}$ is defined by
$C_{n}=\begin{cases}\frac{n!}{(n/2)!(2\pi)^{n}},\,&\textit{if }\,n\,\,\text{is even},\\ 0,\,&\textit{if }\,n\,\,\text{is odd}.\end{cases}$
Proof.
Recall that
$\mathcal{M}_{F}(t)=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+it}}=\frac{-i}{2\pi}\left(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)}{p^{1/2+it}}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-it}}\right).$
Set $x=T^{\delta/3}$ for a suitably small $\delta>0$. Thus we have that the general term of
$\mathcal{M}_{F}(t)^{n}=\frac{(-i)^{n}}{(2\pi)^{n}}\left(\sum_{p\leq T^{\delta}}\frac{A_{F}(1,p)}{p^{1/2+it}}-\sum_{p\leq T^{\delta}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-it}}\right)^{n}$
has the form
$\frac{A_{F}(1,p_{1})^{m(p_{1})}\overline{A_{F}(1,p_{1})^{n(p_{1})}}}{p_{1}^{m(p_{1})(1/2+it)}(-1)^{n(p_{1})}p_{1}^{n(p_{1})(1/2-it)}}\times\cdots\times\frac{A_{F}(1,p_{r})^{m(p_{r})}\overline{A_{F}(1,p_{r})^{n(p_{r})}}}{p_{r}^{m(p_{r})(1/2+it)}(-1)^{n(p_{r})}p_{r}^{n(p_{r})(1/2-it)}},$ (4.4)
where $p_{1}<p_{2}<\cdots<p_{r}\leq T^{\delta}$, $m(p_{j})+n(p_{j})\geq 1$, and $\sum_{j=1}^{r}(m(p_{j})+n(p_{j}))=n$.
Now we discuss the contribution from the general term (4.4) in the following three cases.
[…]
$\begin{split}A_{F}(1,p_{j_{0}})^{m(p_{j_{0}})}=\sum_{k,\ell\geq 0\atop k+\ell\leq m(p_{j_{0}})}a_{k,\ell}(m(p_{j_{0}}))A_{F}(p_{j_{0}}^{k},p_{j_{0}}^{\ell})\end{split}$
and
$\begin{split}\overline{A_{F}(1,p_{j_{0}})^{n(p_{j_{0}})}}=\sum_{k,\ell\geq 0\atop k+\ell\leq n(p_{j_{0}})}a_{k,\ell}(n(p_{j_{0}}))\overline{A_{F}(p_{j_{0}}^{k},p_{j_{0}}^{\ell})}.\end{split}$
Then we can observe that there does not exist $(k,\ell)$ such that $a_{k,\ell}(m(p_{j_{0}}))\neq 0$ and $a_{k,\ell}(n(p_{j_{0}}))\neq 0$. Otherwise, by Lemma 2.1, there would exist $(k,\ell)\in E_{m(p_{j_{0}})}\cap E_{n(p_{j_{0}})}$ and integers $m_{h},\,n_{h}\geq 0$, $h=1,2,3$, such that
$\begin{split}m(p_{j_{0}})&=m_{1}+m_{2}+m_{3},\\ n(p_{j_{0}})&=n_{1}+n_{2}+n_{3},\\ k&=-m_{1}+m_{2}=-n_{1}+n_{2},\\ \ell&=-m_{2}+m_{3}=-n_{2}+n_{3}.\end{split}$
There equations would further lead to
[…]
which contradicts the assumption for Case (I).
With the above observation and the Hecke relations, the numerator of this general term (4.4) can be written as a sum whose terms are all of the form $A_{F}(m_{1},n_{1})\overline{A_{F}(m_{2},n_{2})}$ where $(m_{1},n_{1})\neq(m_{2},n_{2})$. Since for these terms $\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}=0$, we see from Theorem 2.2 and the estimate (1.8) that the contribution from the general term (4.4) to the moment (4.2) is
$\begin{split}&\ll_{n}\frac{1}{\mathcal{H}}\sum_{p_{1}<p_{2}<\cdots<p_{r}\leq T^{\delta}}p_{1}^{-\frac{1}{2}(m(p_{1})+n(p_{1}))}\cdots p_{r}^{-\frac{1}{2}(m(p_{r})+n(p_{r}))}\\ &\qquad\qquad\qquad\qquad\qquad\times\left((T^{1+2n\delta})^{\varepsilon}\big{(}T^{1+n\delta}+T^{3}+T^{1+2n\delta\vartheta}M^{2}\big{)}\right)\\ &\ll_{n}\frac{1}{\mathcal{H}}\bigg{(}\sum_{p\leq T^{\delta}}p^{-\frac{1}{2}}\bigg{)}^{r}\big{(}T^{1+\varepsilon+n\delta(1+2\varepsilon)}+T^{3+\varepsilon+2n\delta\varepsilon}+T^{1+\varepsilon+2n\delta(\vartheta+\varepsilon)}M^{2}\big{)}\\ &\ll_{n}T^{-2+\varepsilon+n\delta(3/2+2\varepsilon)}M^{-2}+T^{\varepsilon+n\delta(1/2+2\varepsilon)}M^{-2}+T^{-2+\varepsilon+n\delta(1/2+2\vartheta+2\varepsilon)},\end{split}$
which is negligible upon taking $0<\delta<\frac{1}{n}$.
[…]
$\sum_{p\leq x}\frac{1}{p}=\log\log x+O(1),$ (4.5)
we have the contribution from this general term to the moment (4.2) is at most
$\begin{split}&\ll_{n}\sum_{p_{1}<p_{2}<\cdots<p_{r}\leq T^{\delta}}p_{1}^{-\frac{1}{2}(m(p_{1})+n(p_{1}))}\cdots p_{r}^{-\frac{1}{2}(m(p_{r})+n(p_{r}))}\\ &\ll_{n}\bigg{(}\sum_{p\leq T^{\delta}}p^{-1}\bigg{)}^{r}\ll_{n}(\log\log T)^{\lfloor\frac{n-1}{2}\rfloor}.\end{split}$
It remains to discuss the following: In the general term (4.4), $m(p_{j})=n(p_{j})=1$ for all $j$. Clearly $n$ must be even, say $n=2m$ and so $r=m$.
Case (III): $m(p_{j})=n(p_{j})=1$ for all $j$. The contribution from these terms to (4.2) is
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\binom{2m}{m}(m!)(m!)\frac{(-1)^{m}(-i)^{2m}}{(2\pi)^{2m}}\\ &\qquad\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{A_{F}(1,p_{1})\overline{A_{F}(1,p_{1})}\cdots A_{F}(1,p_{m})\overline{A_{F}(1,p_{m})}}{p_{1}\cdots p_{m}}.\end{split}$ (4.6)
By the Hecke relations, we have
$\begin{split}&A_{F}(1,p_{1})\overline{A_{F}(1,p_{1})}\cdots A_{F}(1,p_{m})\overline{A_{F}(1,p_{m})}\\ &\quad=A_{F}(1,p_{1}\cdots p_{m})\overline{A_{F}(1,p_{1}\cdots p_{m})}.\end{split}$
Applying Theorem 2.2 and the estimate (4.5), we have (4.6) equals
$\begin{split}&\frac{1}{\mathcal{H}}\frac{(2m)!}{(2\pi)^{2m}}\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{1}{p_{1}\cdots p_{m}}\cdot\left(\mathcal{H}+O\left(T^{1+\varepsilon}(p_{1}\cdots p_{m})+T^{3+\varepsilon}+T^{1+\varepsilon}M^{2}(p_{1}\cdots p_{m})^{2\vartheta}\right)\right)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}+O\bigg{(}T^{-2+\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}1\bigg{)}\\ &+O\bigg{(}T^{\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}\bigg{)}+O\bigg{(}T^{-2+\varepsilon}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{(p_{1}\cdots p_{m})^{1-2\vartheta}}\bigg{)}\\ \\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T+O(1))^{m}+O\big{(}T^{m\delta-2+\varepsilon}M^{-2}+T^{\varepsilon}M^{-2}(\log\log T)^{m}+T^{2m\delta\vartheta-2+\varepsilon}(\log\log T)^{m}\big{)}\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T)^{m}+O_{m}\big{(}(\log\log T)^{m-1}\big{)}.\end{split}$
Now the asymptotic formula (4.2) follows from Cases (I)–(III).
Next we are ready to prove (4.3). Recall that $\sigma_{x}=\sigma_{x,F}$ depending on $F$. By Lemma 4.2,
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}\\ \ll&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}\right|^{2n}+\sum_{k=2}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}\right|^{2n}\\ &+\sum_{k=1}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{2n}\\ &+\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}\big{(}\log(|t|+M_{F})\big{)}^{2n}+\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}.\end{split}$ (4.7)
For the first two terms in (4.7), we have
$\begin{split}&\left|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}\right|^{2n}\ll_{n}\left(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2-it}\log p}\right)^{2n},\end{split}$
and for $k\geq 2$,
$\begin{split}\left|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}\right|^{2n}\ll_{n}\left(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})}{p^{k(1/2-it)}\log p}\right)^{2n}.\end{split}$ (4.8)
For (4.8), we note that
$\begin{split}C_{F}(p^{k})=\alpha_{F,1}(p)^{k}+\alpha_{F,2}(p)^{k}+\alpha_{F,3}(p)^{k}\end{split}$
is a symmetric polynomial in $\alpha_{F,j}(p)$, since (1.5) says that $1$, $A_{F}(1,p)$, $A_{F}(p,1)=\overline{A_{F}(1,p)}$ are the elementary symmetric polynomials in $\alpha_{F,j}(p)$, the fundamental theorem of symmetric polynomials (see [15, pg. 21, Remark]) implies that the symmetric polynomial $C_{F}(p^{k})=\alpha_{F,1}(p)^{k}+\alpha_{F,2}(p)^{k}+\alpha_{F,3}(p)^{k}$ is a polynomial in $1$, $A_{F}(1,p)$ and $\overline{A_{F}(1,p)}$. Then we can use the fact $|\Lambda_{x}(p)-\Lambda(p)|=O\left(\frac{\log^{3}p}{\log^{2}x}\right)$ and the similar argument as Case (I)–(III) to see that the contributions from first two terms in (4.7) to (4.3) is of $O_{t,n}(1)$. The contributions from last two terms in (4.7) is of $O_{t,n}(1)$ by Cauchy–Schwarz inequality, the assumption (1.7), Corollary 2.3 and Lemma 4.1. For the third term in (4.7), it follows from Cauchy–Schwarz inequality that
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{2n}\\ &\leq\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{4n}x^{4nk(\sigma_{x}-1/2)}\right)^{1/2}\\ &\times\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{4n}\right)^{1/2},\end{split}$ (4.9)
where $1\leq k\leq 7$. By Hölder’s inequality with the exponents $4n/(4n-1)$ and $4n$,
$\begin{split}&\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{4n}\\ \leq&\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\mathrm{d}\sigma\right)^{4n-1}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}\mathrm{d}\sigma\right)\\ =&\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}\mathrm{d}\sigma.\end{split}$
Note that
$\begin{split}&\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}\\ &\ll_{n}\left(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}+\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-it)}}\right)^{4n}\\ &\quad+\left(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-it)}}\right)^{4n}.\end{split}$
By using the same argument as in proving (4.2), for $1\leq k\leq 7$, we get
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}&\ll_{t,n}(\log x)^{8n}.\end{split}$
Thus,
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{4n}\\ \ll_{t,n}\,&\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}(\log x)^{8n}\mathrm{d}\sigma\ll_{t,n}(\log x)^{4n}.\end{split}$
Inserting this bound to (4.9) and using Lemma 4.1, we have the third term in (4.7) is of $O_{t,n}(1)$. Thus we complete the proof of this lemma. ∎
Proof of Theorem 1.1. By the binomial theorem, we have
$\begin{split}S_{F}(t)^{n}=\mathcal{M}_{F}(t)^{n}+O_{n}\left(\sum_{\ell=1}^{n}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\right),\end{split}$ (4.10)
where $\mathcal{M}_{F}(t)$ and $\mathcal{R}_{F}(t)$ are defined in (4.1). For $1\leq\ell\leq n$, we apply the generalized Hölder’s inequality, Corollary 2.3 and Lemma 4.3 to get
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\\ \ll&\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\right)^{\frac{1}{2}}\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{2n}\right)^{\frac{n-\ell}{2n}}\\ &\times\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}\right)^{\frac{\ell}{2n}}\\ \ll&{}_{t,n}\,(\log\log T)^{(n-\ell)/2}\ll_{t,n}(\log\log T)^{(n-1)/2}.\end{split}$
Here we use the facts that the positivity of $\mathcal{N}_{F}$ and $h_{T,M}(\mu_{F})\geq 0$ for all $\rm GL_{3}$ Hecke–Maass cusp forms $F$. Then (1.10) follows from (4.2), (4.10) and the above estimate.
For the second assertion of Theorem 1.1, note that the $n$-th moment of $\mu_{TM}$ is
$\int_{\mathbb{R}}\xi^{n}\mathrm{d}\mu_{TM}(\xi)=\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\frac{S_{F}(t)}{\sqrt{\log\log T}}\right)^{n}\right)\bigg{/}\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\right).$
Using (1.10) and Corollary 2.3, we obtain that for all $n\geq 0$,
$\lim_{T\rightarrow\infty}\int_{\mathbb{R}}\xi^{n}\mathrm{d}\mu_{TM}(\xi)=C_{n}=\sqrt{\pi}\int_{\mathbb{R}}\xi^{n}\exp(-\pi^{2}\xi^{2})\mathrm{d}\xi.$
Then by the theory of moments in probability theory (see, for example, [2, Lemma 30.2]), we complete the proof of Theorem 1.1.

### v2

@@
-On an unconditional $\rm GL_{3}$ analog of Selberg’s result
+On an unconditional $\mathrm{GL}_{3}$ analog of Selberg’s result
 Abstract.
-Let $S_{F}(t)=\pi^{-1}\arg L(1/2+it,F)$, where $F$ is a Hecke–Maass cusp form for $\rm SL_{3}(\mathbb{Z})$ in the generic position with the spectral parameter $\nu_{F}=\big{(}\nu_{F,1},\nu_{F,2},\nu_{F,3}\big{)}$ and the Langlands parameter $\mu_{F}=\big{(}\mu_{F,1},\mu_{F,2},\mu_{F,3}\big{)}$.
-In this paper, we establish an unconditional asymptotic formula for the moments of $S_{F}(t)$.
-Previouly, such a formula was only known under the Generalized Riemann Hypothesis.
-The key ingredient is a weighted zero-density estimate in the spectral aspect for $L(s,F)$ which was recently proved by the authors in [18].
+Let $F$ be a Hecke–Maass cusp form for $\mathrm{SL}_{3}(\mathbb{Z})$ with the Langlands parameter $\mu_{F}=\big(\mu_{F,1},\mu_{F,2},\mu_{F,3}\big)$ and the associated $L$-function $L(s,F)$.
+Define $S_{F}(t)=\pi^{-1}\arg L(1/2+\mathrm{i}t,F)$.
+When $\mu_{F}$ is in generic position, we establish an unconditional asymptotic formula for the moments of $S_{F}(t)$.
+Previously, such a formula was only known to hold under the Generalized Riemann Hypothesis.
+The key ingredient is a weighted zero-density estimate in the spectral aspect for $L(s,F)$, which has recently been proved by Sun and Wang in arXiv:2412.02416.
 Key words and phrases:
-Moments, $\rm GL_{3}$ Hecke–Maass cusp forms, Selberg’s limit theorem.
+Moments, $\mathrm{GL}_{3}$ Hecke–Maass cusp forms, Selberg’s limit theorem.
 2010 Mathematics Subject Classification:
 11F12, 11F66, 11F72.
-Q. Sun was partially supported by the National Natural Science Foundation of China (Grant Nos. 12471005 and 12031008) and the Natural Science Foundation of Shandong Province (Grant No. ZR2023MA003).
+†  Corresponding author
 1. Introduction
-It is a well-established fact that the function $S(t)=\frac{1}{\pi}\arg\zeta\left(\frac{1}{2}+it\right)$ is associated with the number of non-trivial zeros of the Riemann zeta function $\zeta(s)$ up to the height $t>0$.
-The argument is defined by continuous variation along the horizontal segment from $\infty+it$ to $1/2+it$, commencing at $\infty+it$ with the value 00.
-The intricate behavior of $S(t)$ represents a highly intriguing topic and has been explored by numerous number theorists (see, e.g., [1, 6, 12, 22]).
-In 1940s, Selberg [19, 20] examined the distribution of $S(t)$ and demonstrated that for any $n\in\mathbb{N}$,
-$\frac{1}{T}\int_{T}^{2T}S(t)^{2n}\mathrm{d}t=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log T)^{n}+O\big{(}(\log\log T)^{n-1/2}\big{)},$ (1.1)
-which implies that, in an average sense, $|S(t)|$ has an order of magnitude $\sqrt{\log\log T}$.
-Additionally, for the Dirichlet $L$-function $L(s,\chi)$ with primitive $\chi$, Selberg [21] studied the behavior of $S(t,\chi)=\frac{1}{\pi}\arg L\big{(}\frac{1}{2}+it,\chi\big{)}$ and obtained, for prime modulus $q$ and any fixed $t>0$,
-$\frac{1}{q}\,\,\sideset{}{{}^{\star}}{\sum}_{\chi\bmod q}S(t,\chi)^{2n}=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log q)^{n}+O\big{(}(\log\log q)^{n-1/2}\big{)},$ (1.2)
-where $\star$ indicates that the summation runs over the primitive characters $\chi(\bmod\,q)$.
+Let $L(s,F)$ be an $L$-function and define
+$\begin{split}S_{F}(t)=\frac{1}{\pi}\arg L\Big(\frac{1}{2}+\mathrm{i}t,F\Big).\end{split}$
+Here, the argument is defined by continuous variation along the horizontal segment from $\infty+\mathrm{i}t$ to $1/2+\mathrm{i}t$, commencing at $\infty+\mathrm{i}t$ with the value 0.
+It is very interesting to study the function $S_{F}(t)$ in number theory.
+For example, in the case of the Riemann zeta function, the classical Riemann–von Mangoldt formula states that the number of zeros $\rho=\beta+\mathrm{i}\gamma$ of $\zeta(s)$ (counted with multiplicity) in the rectangle $\{(\beta,\gamma)\mid 0<\beta<1,0<\gamma\leq t\}$ is given by
+$\frac{t}{2\pi}\log\frac{t}{2\pi\mathrm{e}}+\frac{7}{8}+\frac{1}{\pi}\arg\zeta(1/2+\mathrm{i}t)+O\Big(\frac{1}{t}\Big),$
+which implies that the function $S_{\zeta}(t)=\pi^{-1}\arg\zeta(1/2+\mathrm{i}t)$ is closely related to the nontrivial zeros of $\zeta(s)$.
+As such, the fascinating behavior of $S_{\zeta}(t)$ has been intensively studied by many number theorists (see, e.g., [1, 24, 5, 8] and the references therein).
+On the other hand, the function $S_{F}(t)$ has a close connection with random matrix theory (see [2]).
+In fact, the behavior of $S_{F}(t)$ is analogous to the extreme value statistics of the characteristic polynomials of random matrices.
+To understand more about $S_{F}(t)$, an effective way is to calculate and analyze the various moments of $S_{F}(t)$.
+In the 1940s, Selberg [21, 22] examined the distribution of $S_{\zeta}(t)$ and demonstrated that for any $n\in\mathbb{N}$,
+$\frac{1}{T}\int_{T}^{2T}S_{\zeta}(t)^{2n}\mathrm{d}t=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log T)^{n}+O\big((\log\log T)^{n-1/2}\big),$ (1.1)
+which implies that, in an average sense, $|S_{\zeta}(t)|$ has an order of magnitude $\sqrt{\log\log T}$.
+Additionally, for the Dirichlet $L$-function $L(s,\chi)$ with primitive $\chi$, Selberg [23] studied the behavior of $S_{\chi}(t)=\frac{1}{\pi}\arg L\big(\frac{1}{2}+\mathrm{i}t,\chi\big)$ and obtained, for prime modulus $q$ and any fixed $t>0$,
+$\frac{1}{q}\,\,\sideset{}{{}^{\star}}{\sum}_{\chi\bmod q}S_{\chi}(t)^{2n}=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log q)^{n}+O\big((\log\log q)^{n-1/2}\big),$ (1.2)
+where $\star$ indicates that the summation runs over the primitive characters $\chi$ modulo $q$.
 It should be noted that the counterparts of (1.1) and (1.2) for the odd-order moments can also be derived by making minor adjustments to Selberg’s approach.
-For the $\rm GL_{2}$ family of $L$-functions, Hejhal–Luo [8] first considered the generalization of (1.2) to the case of eigenfunctions $\psi$ of the hyperbolic Laplacian $\Delta$ for the full modular group.
+For the $\mathrm{GL}_{2}$ family of $L$-functions, Hejhal and Luo [11] first considered the generalization of (1.2) to the case of eigenfunctions $\psi$ of the hyperbolic Laplacian $\Delta$ for the full modular group.
 More precisely, let $L(s,\psi)$ be the standard Hecke–Maass $L$-function associated to $\psi$ and define
-$S(t,\psi)=\frac{1}{\pi}\arg L\big{(}\frac{1}{2}+it,\psi\big{)}.$
-They established the analogue of (1.1) when $\psi$ is an Eisenstain series.
-And when $\psi$ is the Hecke–Maass cusp form $u_{j}$ of $\rm SL_{2}(\mathbb{Z})$ with the first Fourier coefficient $\nu_{j}(1)$ and Laplacian eigenvalue $\lambda_{j}=\frac{1}{4}+t_{j}^{2}$ ($t_{j}\geq 0$), under Generalized Riemann Hypothesis (GRH), they showed that for fixed $t,\beta>0$ and any large parameter $T>0$,
-$\begin{split}\sum_{j=1}^{\infty}e^{-\beta(t_{j}-T)^{2}}|\nu_{j}(1)|^{2}S(t,u_{j})^{2n}=\frac{2\pi^{-3/2}}{\sqrt{\beta}}\frac{(2n)!}{n!(2\pi)^{2n}}T(\log\log T)^{n}+O_{t,n,\beta}\big{(}T(\log\log T)^{n-1/2}\big{)}.\end{split}$
-Recently, Sun–Wang [17] reestablished this asymptotic formula by means of the zero-density estimate of weighted type instead of assuming GRH.
-Moreover, for the holomorphic case, Liu–Shim [14] and Sun–Wang [16] successively obtained unconditional analogues in the weight aspect and in the level aspect, respectively.
-For the $\rm GL_{3}$ family of $L$-functions, Liu–Liu [13] explored the average behavior of the analogue function
-$S_{F}(t):=\frac{1}{\pi}\arg L(1/2+it,F)$
-defined for Hecke–Maass cusp forms $F$ for $\rm SL_{3}(\mathbb{Z})$.
-To state the result, we need to introduce some notations.
-Let $F$ run over an orthogonal basis of Hecke–Maass cusp forms for $\rm SL_{3}(\mathbb{Z})$.
-For each $F$, we let $\mu_{F}=\big{(}\mu_{F,1},\mu_{F,2},\mu_{F,3}\big{)}$ be its Langlands parameter and have that $\mu_{F}$ is a point in the region
+$S_{\psi}(t)=\frac{1}{\pi}\arg L\Big(\frac{1}{2}+\mathrm{i}t,\psi\Big).$
+They established the analogue of (1.1) when $\psi$ is an Eisenstein series.
+In addition, when $\psi$ is the Hecke–Maass cusp form $u_{j}$ of $\mathrm{SL}_{2}(\mathbb{Z})$ with the first Fourier coefficient $\nu_{j}(1)$ and Laplacian eigenvalue $\lambda_{j}=1/4+t_{j}^{2}$ ($t_{j}\geq 0$), under the Generalized Riemann Hypothesis (GRH), they showed that for any fixed $t,\beta>0$ and any large positive parameter $T$,
+$\begin{split}\sum_{j=1}^{\infty}\mathrm{e}^{-\beta(t_{j}-T)^{2}}|\nu_{j}(1)|^{2}S_{u_{j}}(t)^{2n}=\frac{2\pi^{-3/2}}{\sqrt{\beta}}\frac{(2n)!}{n!(2\pi)^{2n}}T(\log\log T)^{n}+O_{t,n,\beta}\big(T(\log\log T)^{n-1/2}\big).\end{split}$
+Recently, Sun and Wang [19] reestablished this asymptotic formula by means of a weighted zero-density estimate instead of assuming the GRH.
+Moreover, for the holomorphic case, Liu and Shim [16] and Sun et al.
+[18] successively obtained unconditional analogues in the weight aspect and in the level aspect, respectively.
+For the $\mathrm{GL}_{3}$ family of $L$-functions, Liu and Liu [15] explored the average behavior of the analogue function $S_{F}(t)$, where $F$ is a Hecke–Maass cusp form for $\mathrm{SL}_{3}(\mathbb{Z})$.
+To state their results, we need to introduce some notations.
+For a cusp form $F$ in an orthogonal basis of Hecke–Maass cusp forms for $\mathrm{SL}_{3}(\mathbb{Z})$, we denote its Langlands parameter by $\mu_{F}=\big(\mu_{F,1},\mu_{F,2},\mu_{F,3}\big)$, where $\mu_{F}$ is a point in the region
 $\begin{split}\Lambda_{1/2}^{\prime}=\Bigg\{(\mu_{F,1},\mu_{F,2},\mu_{F,3})\in\mathbb{C}^{3},\,\,\,\begin{aligned} &\mu_{F,1}+\mu_{F,2}+\mu_{F,3}=0,\,\,|\operatorname{Re}(\mu_{F,j})|\leq\frac{1}{2},\,j=1,2,3,\\ &\{-\mu_{F,1},-\mu_{F,2},-\mu_{F,3}\}=\{\overline{\mu_{F,1}},\overline{\mu_{F,2}},\overline{\mu_{F,3}}\}\end{aligned}\Bigg\}\end{split}$
-in the Lie algebra $\mathfrak{sl}_{3}(\mathbb{C})$. Also, the cusp form $F$ has the spectral parameter $\nu_{F}=\big{(}\nu_{F,1},\nu_{F,2},\nu_{F,3}\big{)}$, where
+in the Lie algebra $\mathfrak{sl}_{3}(\mathbb{C})$. Denote by $\nu_{F}=\big(\nu_{F,1},\nu_{F,2},\nu_{F,3}\big)$ the spectral parameter of $F$, where
 $\begin{split}\nu_{F,1}:=\frac{1}{3}&\big(\mu_{F,1}-\mu_{F,2}\big),\quad\nu_{F,2}:=\frac{1}{3}\big(\mu_{F,2}-\mu_{F,3}\big),\\ \nu_{F,3}:=&-\nu_{F,1}-\nu_{F,2}=\frac{1}{3}\big(\mu_{F,3}-\mu_{F,1}\big).\end{split}$ (1.3)
-And we have
+Then we have
 $\mu_{F,1}=2\nu_{F,1}+\nu_{F,2},\quad\mu_{F,2}=\nu_{F,2}-\nu_{F,1},\quad\mu_{F,3}=-\nu_{F,1}-2\nu_{F,2}.$ (1.4)
 We will use the coordinates $\mu$ and $\nu$ simultaneously.
-Note that $\big{(}\nu_{F,1}-\frac{1}{3},\nu_{F,2}-\frac{1}{3}\big{)}$ is the type of $F$ as defined in Goldfeld’s book [7].
-Assume that $F$ has the normalized Fourier coefficients $A_{F}(m,n)$ with $A_{F}(1,1)=1$.
-The $L$-function associated with $F$ is defined by
+Note that $\big(\nu_{F,1}-1/3,\nu_{F,2}-1/3\big)$ is the type of $F$ as defined in Goldfeld’s book [10].
+Assume that $F$ is normalized so that the $(1,1)$-th Fourier coefficient $A_{F}(1,1)=1$.
+The $L$-function associated to $F$ is defined by
 $L(s,F)=\sum_{n=1}^{\infty}\frac{A_{F}(1,n)}{n^{s}},\quad\,\operatorname{Re}(s)>1,$
-and has the Euler product
-$\begin{split}L(s,F)&=\prod_{p}\left(1-\frac{A_{F}(1,p)}{p^{s}}+\frac{A_{F}(p,1)}{p^{2s}}-\frac{1}{p^{3s}}\right)^{-1}\\ &=\prod_{p}\big{(}1-\alpha_{F,1}(p)p^{-s}\big{)}^{-1}\big{(}1-\alpha_{F,2}(p)p^{-s}\big{)}^{-1}\big{(}1-\alpha_{F,3}(p)p^{-s}\big{)}^{-1}.\end{split}$
+which has the Euler product
+$\begin{split}L(s,F)&=\prod_{p}\bigg(1-\frac{A_{F}(1,p)}{p^{s}}+\frac{A_{F}(p,1)}{p^{2s}}-\frac{1}{p^{3s}}\bigg)^{-1}\\ &=\prod_{p}\big(1-\alpha_{F,1}(p)p^{-s}\big)^{-1}\big(1-\alpha_{F,2}(p)p^{-s}\big)^{-1}\big(1-\alpha_{F,3}(p)p^{-s}\big)^{-1}.\end{split}$
 Here we have
-$\begin{split}\begin{cases}\alpha_{F,1}(p)\alpha_{F,2}(p)\alpha_{F,3}(p)=1\\ \alpha_{F,1}(p)+\alpha_{F,2}(p)+\alpha_{F,3}(p)=A_{F}(1,p)\\ \alpha_{F,1}(p)\alpha_{F,2}(p)+\alpha_{F,2}(p)\alpha_{F,3}(p)+\alpha_{F,3}(p)\alpha_{F,1}(p)=A_{F}(p,1),\end{cases}\end{split}$ (1.5)
-and by Kim–Sarnak [10] that
+$\begin{split}\begin{cases}\alpha_{F,1}(p)\alpha_{F,2}(p)\alpha_{F,3}(p)=1,\\ \alpha_{F,1}(p)+\alpha_{F,2}(p)+\alpha_{F,3}(p)=A_{F}(1,p),\\ \alpha_{F,1}(p)\alpha_{F,2}(p)+\alpha_{F,2}(p)\alpha_{F,3}(p)+\alpha_{F,3}(p)\alpha_{F,1}(p)=A_{F}(p,1),\end{cases}\end{split}$ (1.5)
+and by Kim and Sarnak [13],
 $\begin{split}|\alpha_{F,j}(p)|\leq p^{\frac{5}{14}},\quad\text{and}\quad|\operatorname{Re}(\mu_{F,j})|\leq\frac{5}{14},\quad\text{for}\,\,j=1,2,3.\end{split}$ (1.6)
 Note that $L(s,F)$ has an analytic continuation to the whole complex plane and satisfies the functional equation
-$\Lambda(s,F)=L(s,F)\prod_{j=1}^{3}\Gamma_{\mathbb{R}}(s-\mu_{F,j})=L(1-s,\widetilde{F})\prod_{j=1}^{3}\Gamma_{\mathbb{R}}(1-s+\mu_{F,j})=\Lambda(1-s,\widetilde{F}),$
+[…]
 where $\Gamma_{\mathbb{R}}(s)=\pi^{-s/2}\Gamma(s/2)$ and the dual form of $F$ is denoted by $\widetilde{F}$ with the Langlands parameter $(-\mu_{F,1},-\mu_{F,2},-\mu_{F,3})$, the spectral parameter $\big(\nu_{F,2},\nu_{F,1},\nu_{F,3}\big)$ and the coefficients $A_{\widetilde{F}}(m,n)=\overline{A_{F}(m,n)}=A_{F}(n,m)$.
-Let $\mu_{0}=(\mu_{0,1},\mu_{0,2},\mu_{0,3})$ and $\nu_{0}=(\nu_{0,1},\nu_{0,2},\nu_{0,3})$, and satisfy the corresponding relations (1.3) and (1.4). We also assume
+Let $\mu_{0}=(\mu_{0,1},\mu_{0,2},\mu_{0,3})$ and $\nu_{0}=(\nu_{0,1},\nu_{0,2},\nu_{0,3})$, which satisfy the corresponding relations (1.3) and (1.4).
+We also assume that $\mu_{0}$ is in generic position, i.e.,
 $|\mu_{0,j}|\asymp|\nu_{0,j}|\asymp T:=\|\mu_{0}\|,\quad 1\leq j\leq 3,$ (1.8)
@@
 $\|\mu_{0}\|=\sqrt{|\mu_{0,1}|^{2}+|\mu_{0,2}|^{2}+|\mu_{0,3}|^{2}}.$
-Throughout this paper, we let $M\!=\!T^{1-\eta}$ for some very small $20<\eta<1/$ 2.
+Throughout this paper, we let $M\!=\!T^{1-\eta}$ for some sufficiently small $0<\eta<1/2$.
 Define a test function $h_{T,M}(\mu)$ (depending on $\mu_{0}$) for $\mu=(\mu_{1},\mu_{2},\mu_{3})\in\mathbb{C}^{3}$ by
 $h_{T,M}(\mu):=P(\mu)^{2}\bigg(\sum_{w\in\mathcal{W}}\psi\bigg(\frac{w(\mu)-\mu_{0}}{M}\bigg)\bigg)^{2},$
-where $\psi(\mu)=\exp\left(\mu_{1}^{2}+\mu_{2}^{2}+\mu_{3}^{2}\right)$ and
-$P(\mu)=\prod_{0\leq n\leq A}\prod_{j=1}^{3}\frac{\left(\nu_{j}-\frac{1}{3}(1+2n)\right)\left(\nu_{j}+\frac{1}{3}(1+2n)\right)}{|\nu_{0,j}|^{2}}$
+where $\psi(\mu)=\exp\big(\mu_{1}^{2}+\mu_{2}^{2}+\mu_{3}^{2}\big)$ and
+$P(\mu)=\prod_{0\leq n\leq A}\prod_{j=1}^{3}\frac{\big(\nu_{j}-\frac{1}{3}(1+2n)\big)\big(\nu_{j}+\frac{1}{3}(1+2n)\big)}{|\nu_{0,j}|^{2}}$
 for some fixed large $A>0$. Here
-$\mathcal{W}:=\left\{I,\;w_{2}=\left(\begin{smallmatrix}1&&\\ &&-1\\ &1&\end{smallmatrix}\right),\;w_{3}=\left(\begin{smallmatrix}&1&\\ -1&&\\ &&1\end{smallmatrix}\right),\;w_{4}=\left(\begin{smallmatrix}&1&\\ &&1\\ 1&&\end{smallmatrix}\right),\;w_{5}=\left(\begin{smallmatrix}&&1\\ 1&&\\ &1&\end{smallmatrix}\right),\;w_{6}=\left(\begin{smallmatrix}&&1\\ &-1&\\ 1&&\end{smallmatrix}\right)\right\}$
-is the Weyl group of $\rm SL_{3}(\mathbb{R})$.
-The function $h_{T,M}(\mu)$ localizes at a ball of radius $M$ about $w(\mu_{0})$ for each $w\in\mathcal{W}$.
+$\mathcal{W}:=\Big\{I,\;w_{2}=\Big(\begin{smallmatrix}1&&\\ &&-1\\ &1&\end{smallmatrix}\Big),\;w_{3}=\Big(\begin{smallmatrix}&1&\\ -1&&\\ &&1\end{smallmatrix}\Big),\;w_{4}=\Big(\begin{smallmatrix}&1&\\ &&1\\ 1&&\end{smallmatrix}\Big),\;w_{5}=\Big(\begin{smallmatrix}&&1\\ 1&&\\ &1&\end{smallmatrix}\Big),\;w_{6}=\Big(\begin{smallmatrix}&&1\\ &-1&\\ 1&&\end{smallmatrix}\Big)\Big\}$
+is the Weyl group of $\mathrm{SL}_{3}(\mathbb{R})$. The function $h_{T,M}(\mu)$ localizes at a ball of radius $M$ about $w(\mu_{0})$ for each $w\in\mathcal{W}$.
 Moreover, we define the spectral measure on the hyperplane $\mu_{1}+\mu_{2}+\mu_{3}=0$ by
@@
 where
-$\text{spec}(\mu):=\prod_{j=1}^{3}\left(3\nu_{j}\tan\left(\frac{3\pi}{2}\nu_{j}\right)\right)\quad\mbox{and}\quad\mathrm{d}\mu=\mathrm{d}\mu_{1}\mathrm{d}\mu_{2}=\mathrm{d}\mu_{2}\mathrm{d}\mu_{3}=\mathrm{d}\mu_{1}\mathrm{d}\mu_{3}.$
-By trivial estimate, we have
+$\text{spec}(\mu):=\prod_{j=1}^{3}\bigg(3\nu_{j}\tan\bigg(\frac{3\pi}{2}\nu_{j}\bigg)\bigg)\quad\mbox{and}\quad\mathrm{d}\mu=\mathrm{d}\mu_{1}\mathrm{d}\mu_{2}=\mathrm{d}\mu_{2}\mathrm{d}\mu_{3}=\mathrm{d}\mu_{1}\mathrm{d}\mu_{3}.$
+By a trivial estimate, we have
 $\mathcal{H}:=\frac{1}{192\pi^{5}}\int_{\operatorname{Re}(\mu)=0}h_{T,M}(\mu)\mathrm{d}_{\rm{spec}}\mu\asymp T^{3}M^{2}.$ (1.9)
-Fix any $t>0$ and $n\in\mathbb{N}$, under GRH, Liu and Liu [13] established that, for large positive $T$,
+Fix any $t>0$ and $n\in\mathbb{N}$, under the GRH, Liu and Liu [15] established that, for large positive $T$,
 $\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\nu_{F})}{L(1,\text{Ad}^{2}F)}S_{F}(t)^{2n}=\frac{9(2n)!}{4n!(2\pi)^{2n}}(\log\log T)^{n}+O_{t,n}\big((\log\log T)^{n-1/2}\big),\end{split}$ (1.10)
-where $L(1,\text{Ad}^{2}F)$ denotes the adjoint square $L$-function of $F$.
+where $L(s,\text{Ad}^{2}F)$ denotes the adjoint square $L$-function of $F$.
+Recently, with the automorphic Plancherel density theorem of Matz and Templier, Chen et al.
+[9] extended their result to $\mathrm{GL}_{n}$.
 In this paper, we are devoted to proving the asymptotic formula (1.10) without assuming the GRH.
 We introduce the normalizing factor
-$\mathcal{N}_{F}:=\|F\|^{2}\prod_{j=1}^{3}\cos\left(\frac{3}{2}\pi\nu_{F,j}\right).$
-By Rankin–Selberg theory in combination with Stade’s formula (see e.g., Blomer [3, Section 4] and Li [11, Theorem 2]), it is easy to see that
-$\begin{split}\mathcal{N}_{F}\asymp\mathop{\operatorname{Res}}_{s=1}L(s,F\times\widetilde{F})\ll\|\mu_{F}\|^{\varepsilon},\end{split}$
-where $L(s,F\times\widetilde{F})$ is the Rankin–Selberg square of $F$ and the implied constants depending at most on $\varepsilon$. Our main result is the following theorem.
+$\mathcal{N}_{F}:=\|F\|^{2}\prod_{j=1}^{3}\cos\bigg(\frac{3}{2}\pi\nu_{F,j}\bigg).$
+By Rankin–Selberg theory in combination with Stade’s formula (see e.g., Blomer [4, Section 4] and Li [14, Theorem 2]), it is easy to see that
+$\begin{split}\mathcal{N}_{F}\asymp\mathrm{Res}_{s=1}L(s,F\times\widetilde{F})\ll\|\mu_{F}\|^{\varepsilon},\end{split}$
+[…]
+$\begin{split}L(s,F\times\widetilde{F})&=\zeta(3s)\sum_{m,n\geq 1}\frac{|A_{F}(m,n)|^{2}}{(m^{2}n)^{s}}\\ &=\prod_{p}\prod_{i=1}^{3}\prod_{j=1}^{3}\bigg(1-\alpha_{F,i}(p)\overline{\alpha_{F,j}(p)}p^{-s}\bigg)^{-1},\quad\text{for}\,\,\,\operatorname{Re}(s)>1,\end{split}$
+and the implied constants depend at most on $\varepsilon$. Our main result is the following theorem.
 Theorem 1.1.
-Let $t>0$ be fixed and $T$ be any large parameter, we have
+Let $t>0$ be fixed and $T$ be any large positive parameter. We have
 $\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}S_{F}(t)^{n}=C_{n}(\log\log T)^{n/2}+O_{t,n}\big((\log\log T)^{(n-1)/2}\big),\end{split}$ (1.11)
@@
 $C_{n}=\begin{cases}\frac{n!}{(n/2)!(2\pi)^{n}},\,&\textit{if }\,n\,\,\text{is even},\\ 0,\,&\textit{if }\,n\,\,\text{is odd}.\end{cases}$
-Moreover, we define the following probability measure $\mu_{TM}$ on $\mathbb{R}$ by
-$\begin{split}\mu_{TM}(E):=\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\mathbf{1}_{E}\left(\frac{S_{F}(t)}{\sqrt{\log\log T}}\right)\right)\bigg{/}\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\right),\end{split}$
+Moreover, define the following probability measure $\mu_{TM}$ on $\mathbb{R}$ by
+$\begin{split}\mu_{TM}(E):=\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\mathbf{1}_{E}\bigg(\frac{S_{F}(t)}{\sqrt{\log\log T}}\bigg)\bigg)\bigg/\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg),\end{split}$
 where $\mathbf{1}_{E}$ is the characteristic function of a Borel measurable set $E$ in $\mathbb{R}$. Then for every $a<b$, we have
@@
 Remark 1.
-For $F$ a Hecke–Maass cusp form of $\rm SL_{3}(\mathbb{Z})$, on an average sense, our result indicates that $|S_{F}(t)|$ has order of magnitude $\sqrt{\log\log T}$ under the assumption (1.7).
-In addition, as $T\rightarrow\infty$, the ratios $S_{F}(t)/\sqrt{\log\log T}$ converges in distribution to a normal distribution of mean 00 and variance $(2\pi^{2})^{-1}$.
+For $F$ a Hecke–Maass cusp form of $\mathrm{SL}_{3}(\mathbb{Z})$, our result indicates that, as $T\rightarrow\infty$, the ratios $S_{F}(t)/\sqrt{\log\log T}$ converge in distribution to a normal distribution of mean 0 and variance $(2\pi^{2})^{-1}$.
 Notation Throughout the paper, $\varepsilon$ is an arbitrarily small positive real number, which may be different at each occurrence.
-As usual, $e(z)=\exp(2\pi iz)=e^{2\pi iz}$ and $\delta_{m,n}$ is the Kronecker symbol that detects $m=n$.
+As usual, $e(z)=\exp(2\pi\mathrm{i}z)=\mathrm{e}^{2\pi\mathrm{i}z}$ and $\delta_{m,n}$ is the Kronecker symbol that detects $m=n$.
 We write $f=O(g)$ or $f\ll g$ to mean $|f|\leq cg$ for some unspecified positive constant $c$.
@@
 Structure of this paper.
-This paper is organized as follows.
-In Sect.
-2, we give some preliminary results needed for the proof.
-In Sect.
-3, we establish some lemmas and provide a crucial approximation of $S_{F}(t)$.
-In Sect.
-4, we give the detailed proof of Theorem 1.1.
+The rest of this paper is organized as follows.
+In Section 2, we give some preliminary results needed for the proof.
+In Section 3, we establish some lemmas and provide a crucial approximation of $S_{F}(t)$.
+In Section 4, we give the detailed proof of Theorem 1.1.
+In Section 5, we summarize our main results, highlight the significance of our work, and discuss prospects for future research.
 2. Preliminaries
 2.1. Hecke relations
-Firstly, let’s recall the Hecke relations (see Goldfeld [7, pg. 168]):
-$\begin{split}A_{F}(m_{1}m_{2},n_{1}n_{2})&=A_{F}(m_{1},n_{1})A_{F}(m_{2},n_{2}),\quad\text{if}\,\,(m_{1}n_{1},m_{2}n_{2})=1,\\ A_{F}(n,1)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\left(\frac{m_{1}d_{0}}{d_{1}},\frac{m_{2}d_{1}}{d_{2}}\right),\\ A_{F}(1,n)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\left(\frac{m_{1}d_{2}}{d_{1}},\frac{m_{2}d_{0}}{d_{2}}\right),\\ A_{F}(n_{1},1)A_{F}(1,n_{2})&=\sum_{d|(n_{1},n_{2})}A_{F}\left(\frac{n_{1}}{d},\frac{n_{2}}{d}\right).\end{split}$
-Then we arrive at the following useful result, which was presented by Liu–Liu in [13, Proposition 5.1].
+Firstly, let us recall the following Hecke relations (see Goldfeld [10, pg. 168])
+$\begin{split}A_{F}(m_{1}m_{2},n_{1}n_{2})&=A_{F}(m_{1},n_{1})A_{F}(m_{2},n_{2}),\quad\text{if}\,\,(m_{1}n_{1},m_{2}n_{2})=1,\\ A_{F}(n,1)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\bigg(\frac{m_{1}d_{0}}{d_{1}},\frac{m_{2}d_{1}}{d_{2}}\bigg),\\ A_{F}(1,n)A_{F}(m_{1},m_{2})&=\sum_{d_{0}d_{1}d_{2}=n\atop d_{1}|m_{1},\,d_{2}|m_{2}}A_{F}\bigg(\frac{m_{1}d_{2}}{d_{1}},\frac{m_{2}d_{0}}{d_{2}}\bigg),\\ A_{F}(n_{1},1)A_{F}(1,n_{2})&=\sum_{d|(n_{1},n_{2})}A_{F}\bigg(\frac{n_{1}}{d},\frac{n_{2}}{d}\bigg).\end{split}$
+Then we arrive at the following useful result, which was presented by Liu and Liu in [15, Proposition 5.1].
 Lemma 2.1.
@@
 2.2. A summation formula
-We shall need the following summation formula, which is an application of the $\rm GL_{3}$ Kuznetsov trace formula in the version of Buttcane [4, Theorems 2–4].
+We need the following summation formula, which is an application of the $\mathrm{GL}_{3}$ Kuznetsov trace formula in the version of Buttcane [6, Theorems 2–4].
 Theorem 2.2.
-Let $P=m_{1}n_{1}m_{2}n_{2}\neq 0$. We have for any $\varepsilon>0$ that
-$\begin{split}&\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\overline{A_{F}(m_{1},m_{2})}A_{F}(n_{1},n_{2})\\ &\quad=\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}\mathcal{H}+O\left((TP)^{\varepsilon}\big{(}TP^{1/2}+T^{3}+TM^{2}P^{\vartheta}\big{)}\right),\end{split}$
-where $\vartheta=7/64$ is the best result towards the Ramanujan–Petersson conjecture for Fourier coefficients of $\rm GL_{2}$ Maass cusp forms due to Kim–Sarnak [10].
+Let $P=m_{1}n_{1}m_{2}n_{2}\neq 0$. For any $\varepsilon>0$, we have
+$\begin{split}&\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\overline{A_{F}(m_{1},m_{2})}A_{F}(n_{1},n_{2})\\ &\quad=\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}\mathcal{H}+O\bigg((TP)^{\varepsilon}\big(TP^{1/2}+T^{3}+TM^{2}P^{\vartheta}\big)\bigg),\end{split}$
+where $\vartheta=7/64$, due to Kim and Sarnak [13], is the best record towards the Ramanujan–Petersson conjecture for Fourier coefficients of $\mathrm{GL}_{2}$ Maass cusp forms.
 Proof.
-See Sun–Wang [18, Theorem 7.1]. ∎
-We record here the frequently used estimate
+See Sun and Wang [20, Theorem 7.1]. ∎
+We quote here the frequently used estimate
 $\mathcal{H}=CT^{3}M^{2}+O(T^{2}M^{3}),$
-for some constant $C>0$ (see Buttcane–Zhou [5, pg. 15]). Taking $m_{1}=n_{1}=m_{2}=n_{2}=1$ in Theorem 2.2 immediately gives
+for some constant $C>0$ (see Buttcane and Zhou [7, pg. 15]). Taking $m_{1}=n_{1}=m_{2}=n_{2}=1$ in Theorem 2.2 immediately gives
 Corollary 2.3.
-With the assumptions in Theorem 2.2. We have
+Let the assumptions be as in Theorem 2.2. We have
 $\begin{split}&\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}=CT^{3}M^{2}+O(T^{2}M^{3}).\end{split}$
 2.3. The zero-density estimate
-We will need the following weighted version of zero-density estimate in the spectral aspect for the $L$-functions associated to $\rm GL_{3}$ Hecke–Maass cusp forms which was established by Sun–Wang [18, Theorem 1.1].
+In the context of our proof, the key point lies in the following weighted version of zero-density estimate in the spectral aspect for the $L$-functions associated to $\mathrm{GL}_{3}$ Hecke–Maass cusp forms, which was rigorously established by Sun and Wang in [20, Theorem 1.1].
 Lemma 2.4.
 For $\sigma>0$ and $H>0$, we let
-$\begin{split}N(\sigma,H;F):=\#\big{\{}\rho=1/2+\beta+i\gamma:L(\rho,F)=0,\beta>\sigma,|\gamma|<H\big{\}}\end{split}$
+$\begin{split}N(\sigma,H;F):=\#\big\{\rho=1/2+\beta+\mathrm{i}\gamma:L(\rho,F)=0,\beta>\sigma,|\gamma|<H\big\}\end{split}$
 and
 $\begin{split}N(\sigma,H):=\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}N(\sigma,H;F).\end{split}$
-Let $22/\log T<\sigma<1/$ 2. For some sufficiently small $\delta_{1},\,\theta>0$, we have
+Let $2/\log T<\sigma<1/2$. For some sufficiently small $\delta_{1},\,\theta>0$, we have
 $\begin{split}N(\sigma,H)\ll HT^{-\theta\sigma}\log T,\end{split}$
@@
 3. An approximation of $S_{F}(t)$
-In this section we will prove several lemmas and derive an approximation of $S_{F}(t)$.
-We denote by $\rho=\beta+i\gamma$ a typical zero of $L(s,F)$ inside the critical strip, i.e., $10<\beta<$ 1.
+In this section we prove several lemmas and derive an approximation of $S_{F}(t)$.
+We denote by $\rho=\beta+\mathrm{i}\gamma$ a typical (non-trivial) zero of $L(s,F)$ inside the critical strip, i.e., $0<\beta<1$.
 For $\operatorname{Re}(s)>1$, we have the Dirichlet series expansion for the logarithmic derivative of $L(s,F)$,
-$-\frac{L^{\prime}}{L}(s,F)=\sum_{n=1}^{\infty}\frac{\Lambda(n)C_{F}(n)}{n^{s}},$ (3.1)
+$\frac{L^{\prime}}{L}(s,F)=\sum_{n=1}^{\infty}\frac{\Lambda(n)C_{F}(n)}{n^{s}},$ (3.1)
 where $\Lambda(n)$ denotes the von Mangoldt function, and
@@
 Lemma 3.1.
-Let $x>1$. For $s\neq\rho$, and $s\neq-2m+\mu_{F,j}$, ($m\geq 0$, $j=1,2,3$), we have the following identity
+Let $x>1$. For $s\neq\rho$ and $s\neq-2m+\mu_{F,j}$ ($m\geq 0$, $j=1,2,3$), we have the following identity
 $\begin{split}\frac{L^{\prime}}{L}(s,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{s}}-\frac{1}{\log^{2}x}\sum_{\rho}\frac{x^{\rho-s}(1-x^{\rho-s})^{2}}{(\rho-s)^{3}}\\ &-\frac{1}{\log^{2}x}\sum_{j=1}^{3}\sum_{m=0}^{\infty}\frac{x^{-2m-s+\mu_{F,j}}(1-x^{-2m-s+\mu_{F,j}})^{2}}{(-2m-s+\mu_{F,j})^{3}},\end{split}$
 where
-$\Lambda_{x}(n)=\begin{cases}\Lambda(n),\,&\textit{if }\,\,n\leq x,\\ \Lambda(n)\frac{\log^{2}(x^{3}/n)-2\log^{2}(x^{2}/n)}{2\log^{2}x},\,&\textit{if }\,\,x\leq n\leq x^{2},\\ \Lambda(n)\frac{\log^{2}(x^{3}/n)}{2\log^{2}x},\,&\textit{if }\,\,x^{2}\leq n\leq x^{3},\\ 0,\,&\textit{if }\,\,n\geq x^{3}.\\ \end{cases}$
+[…]
 Proof.
 Recall the discontinuous integral
-$\frac{1}{2\pi i}\int_{(\alpha)}\frac{y^{s}}{s^{3}}\mathrm{d}s=\begin{cases}\frac{\log^{2}y}{2},\,&\textit{if }\,\,y\geq 1,\\ 0,\,&\textit{if }\,\,0<y\leq 1\end{cases}$
+$\frac{1}{2\pi\mathrm{i}}\int_{(\alpha)}\frac{y^{s}}{s^{3}}\mathrm{d}s=\begin{cases}\frac{\log^{2}y}{2},\,&\textit{if }\,\,y\geq 1,\\ 0,\,&\textit{if }\,\,0<y\leq 1\end{cases}$
 for $\alpha>0$. It follows from (3.1) and (3.2) that
-$-\log^{2}x\sum_{n=1}^{\infty}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{s}}=\frac{1}{2\pi i}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v,$
+$\log^{2}x\sum_{n=1}^{\infty}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{s}}=\frac{1}{2\pi\mathrm{i}}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v,$
 where $\alpha=\max\{2,1+\operatorname{Re}(s)\}$.
 By moving the line of integration all way to the left, we pick up the residues at $v=0$, $v=\rho-s$ and $v=-2m-s+\mu_{F,j}$ ($m\geq 0$, $j=1,2,3$), and deduce that
-$\begin{split}&\frac{1}{2\pi i}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v\\ =&\frac{L^{\prime}}{L}(s,F)\log^{2}x+\sum_{\rho}\frac{x^{\rho-s}(1-x^{\rho-s})^{2}}{(\rho-s)^{3}}+\sum_{j=1}^{3}\sum_{m=0}^{\infty}\frac{x^{-2m-s+\mu_{F,j}}(1-x^{-2m-s+\mu_{F,j}})^{2}}{(-2m-s+\mu_{F,j})^{3}}.\end{split}$
-Thus we complete the proof of lemma. ∎
+$\begin{split}&\frac{1}{2\pi\mathrm{i}}\int_{(\alpha)}\frac{x^{v}(1-x^{v})^{2}}{v^{3}}\frac{L^{\prime}}{L}(s+v,F)\mathrm{d}v\\ =&\frac{L^{\prime}}{L}(s,F)\log^{2}x+\sum_{\rho}\frac{x^{\rho-s}(1-x^{\rho-s})^{2}}{(\rho-s)^{3}}+\sum_{j=1}^{3}\sum_{m=0}^{\infty}\frac{x^{-2m-s+\mu_{F,j}}(1-x^{-2m-s+\mu_{F,j}})^{2}}{(-2m-s+\mu_{F,j})^{3}}.\end{split}$
+This completes the proof of the lemma. ∎
 Lemma 3.2.
-For $s=\sigma+it$, $s^{\prime}=\sigma^{\prime}+it^{\prime}$ such that $1/2\leq\sigma,\sigma^{\prime}\leq 2$, $s\neq\rho$ and $s^{\prime}\neq\rho$, we have
-$\operatorname{Im}\left(\frac{L^{\prime}}{L}(s,F)-\frac{L^{\prime}}{L}(s^{\prime},F)\right)=\operatorname{Im}\sum_{\rho}\left(\frac{1}{s-\rho}-\frac{1}{s^{\prime}-\rho}\right)+O(1),$
+For $s=\sigma+\mathrm{i}t$ and $s^{\prime}=\sigma^{\prime}+\mathrm{i}t^{\prime}$ such that $1/2\leq\sigma,\sigma^{\prime}\leq 10$, $s\neq\rho$ and $s^{\prime}\neq\rho$, we have
+$\operatorname{Im}\bigg(\frac{L^{\prime}}{L}(s,F)-\frac{L^{\prime}}{L}(s^{\prime},F)\bigg)=\operatorname{Im}\sum_{\rho}\bigg(\frac{1}{s-\rho}-\frac{1}{s^{\prime}-\rho}\bigg)+O(1),$
 and
@@
 Proof.
-By the Hadamard factorization theorem of the entire function $\Lambda(s,F)$, we have
-$\frac{L^{\prime}}{L}(s,F)=b_{F}+\sum_{\rho}\left(\frac{1}{s-\rho}+\frac{1}{\rho}\right)-\frac{1}{2}\sum_{j=1}^{3}\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{s-\mu_{F,j}}{2}\right)+\frac{3}{2}\log\pi,$
-for some $b_{F}\in\mathbb{C}$ with $\operatorname{Re}(b_{F})=-\operatorname{Re}\sum_{\rho}\frac{1}{\rho}$ (see [9, Proposition 5.7 (3)]). Note that for $z\not\in-\mathbb{N}$,
-$\frac{\Gamma^{\prime}}{\Gamma}(z)=-\gamma+\sum_{k\geq 1}\left(\frac{1}{k}-\frac{1}{k-1+z}\right),$ (3.4)
-where $\gamma$ is the Euler constant. Using (3.4), we get that for $\frac{1}{4}\leq u$,
-$\operatorname{Im}\frac{\Gamma^{\prime}}{\Gamma}(u+iv)=\sum_{k\geq 1}\frac{v}{(k+u-1)^{2}+v^{2}}\ll 1.$
-Combined with the asymptotic formula of the logarithmic derivative of $\Gamma(z)$,
-$\frac{\Gamma^{\prime}}{\Gamma}(z)=\log z-\frac{1}{2z}+O_{\varepsilon}\left(\frac{1}{|z|^{2}}\right),\quad\text{for}\,\,|\arg z|\leq\pi-\varepsilon,$
-we complete the proof of lemma. ∎
+By the Hadamard factorization theorem for the entire function $\Lambda(s,F)$, we have
+$\frac{L^{\prime}}{L}(s,F)=b_{F}+\sum_{\rho}\bigg(\frac{1}{s-\rho}+\frac{1}{\rho}\bigg)-\frac{1}{2}\sum_{j=1}^{3}\frac{\Gamma^{\prime}}{\Gamma}\bigg(\frac{s-\mu_{F,j}}{2}\bigg)+\frac{3}{2}\log\pi,$
+for some $b_{F}\in\mathbb{C}$ with $\operatorname{Re}(b_{F})=-\operatorname{Re}\sum_{\rho}\frac{1}{\rho}$ (see [12, Proposition 5.7 (3)]). Note that for $z\not\in$-$\mathbb{N}$,
+$\frac{\Gamma^{\prime}}{\Gamma}(z)=-\gamma+\sum_{k\geq 1}\bigg(\frac{1}{k}-\frac{1}{k-1+z}\bigg),$ (3.4)
+where $\gamma$ is the Euler constant. Using (3.4), we get that for $\frac{1}{14}\leq u\leq\frac{145}{28}$ and $v\ll 1$,
+$\operatorname{Im}\frac{\Gamma^{\prime}}{\Gamma}(u+\mathrm{i}v)=\sum_{k\geq 1}\frac{v}{(k+u-1)^{2}+v^{2}}\ll 1.$
+Combining with the asymptotic formula of the logarithmic derivative of $\Gamma(z)$,
+$\frac{\Gamma^{\prime}}{\Gamma}(z)=\log z-\frac{1}{2z}+O_{\varepsilon}\bigg(\frac{1}{|z|^{2}}\bigg),\quad\text{for}\,\,|\arg z|\leq\pi-\varepsilon,$
+we complete the proof of the lemma. ∎
 Let $x\geq 4$. We define
-$\sigma_{x}=\sigma_{x,F}=\sigma_{x,F,t}:=\frac{1}{2}+2\max_{\rho}\left\{\left|\beta-\frac{1}{2}\right|,\frac{5}{\log x}\right\},$ (3.5)
-where $\rho=\beta+i\gamma$ runs through the zeros of $L(s,F)$ for which
+$\sigma_{x}=\sigma_{x,F}=\sigma_{x,F,t}:=\frac{1}{2}+2\max_{\rho}\bigg\{\bigg|\beta-\frac{1}{2}\bigg|,\frac{5}{\log x}\bigg\},$ (3.5)
+where $\rho=\beta+\mathrm{i}\gamma$ runs through the zeros of $L(s,F)$ for which
 $|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}.$
@@
 Let $x\geq 4$. For $\sigma\geq\sigma_{x}$, we have
-$\begin{split}\frac{L^{\prime}}{L}(\sigma+it,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}+O\left(x^{1/4-\sigma/2}\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)\\ &+O\left(x^{1/4-\sigma/2}\log(|t|+M_{F})\right),\end{split}$
+$\begin{split}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}+O\bigg(x^{1/4-\sigma/2}\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)\\ &+O\bigg(x^{1/4-\sigma/2}\log(|t|+M_{F})\bigg),\end{split}$
 and
-$\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}=O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\big{(}\log(|t|+M_{F})\big{)}.$
+$\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}=O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\big(\log(|t|+M_{F})\big).$
 Proof.
 By (3.3), we have
-$\operatorname{Re}\frac{L^{\prime}}{L}(\sigma_{x}+it,F)=\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\log(|t|+M_{F})).$ (3.6)
-Moreover, if $\rho=\beta+i\gamma$ is zero of $L(s,F)$, then $(1-\beta)+i\gamma$ is also a zero of $L(s,F)$. Thus we have
-$\begin{split}&\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\\ =&\frac{1}{2}\left(\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+\sum_{\rho}\frac{\sigma_{x}-(1-\beta)}{(\sigma_{x}-(1-\beta))^{2}+(t-\gamma)^{2}}\right)\\ =&\left(\sigma_{x}-\frac{1}{2}\right)\sum_{\rho}\frac{(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})((\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2})}.\end{split}$ (3.7)
+$\operatorname{Re}\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)=\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\log(|t|+M_{F})).$ (3.6)
+Moreover, if $\rho=\beta+\mathrm{i}\gamma$ is zero of $L(s,F)$, then $(1-\beta)+\mathrm{i}\gamma$ is also a zero of $L(s,F)$ by the functional equation (1.7). Thus we have
+$\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}$
+$=$ $\,\frac{1}{2}\bigg(\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+\sum_{\rho}\frac{\sigma_{x}-(1-\beta)}{(\sigma_{x}-(1-\beta))^{2}+(t-\gamma)^{2}}\bigg)$
+$=$ $\,\bigg(\sigma_{x}-\frac{1}{2}\bigg)\sum_{\rho}\frac{(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})((\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2})}.$ (3.7)
 If $|\beta-1/2|\leq\frac{\sigma_{x}-1/2}{2}$, then
@@
 $(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}\geq\frac{1}{4}\big((\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2}\big).$ (3.8)
-If $|\beta-1/2|>\frac{\sigma_{x}-1/2}{2}$, then by (3.5) and (3) we have
+If $|\beta-1/2|>\frac{\sigma_{x}-1/2}{2}$, then by (3.5) we have
 $|t-\gamma|>\frac{x^{3|\beta-1/2|}}{\log x}>3|\beta-1/2|.$
 Thus,
-$\begin{split}&(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}\\ &=\big{(}(\sigma_{x}-1/2)^{2}+(\beta-1/2)^{2}\big{)}+(t-\gamma)^{2}-2(\beta-1/2)^{2}\\ &\geq\frac{1}{2}\big{(}(\sigma_{x}-\beta)^{2}+(\sigma_{x}-1+\beta)^{2}\big{)}+\frac{7}{9}(t-\gamma)^{2}\\ &\geq\frac{1}{4}\big{(}(\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2}\big{)}.\end{split}$ (3.9)
-Combining (3.8), (3.9) and (3.7), we have
+$(\sigma_{x}-1/2)^{2}-(\beta-1/2)^{2}+(t-\gamma)^{2}$
+$=\big((\sigma_{x}-1/2)^{2}+(\beta-1/2)^{2}\big)+(t-\gamma)^{2}-2(\beta-1/2)^{2}$
+$\geq\frac{1}{2}\big((\sigma_{x}-\beta)^{2}+(\sigma_{x}-1+\beta)^{2}\big)+\frac{7}{9}(t-\gamma)^{2}$
+$\geq\frac{1}{4}\big((\sigma_{x}-1+\beta)^{2}+(t-\gamma)^{2}\big).$ (3.9)
+Combining (3.8), (3) and (3), we get
 $\sum_{\rho}\frac{\sigma_{x}-\beta}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}\geq\frac{1}{4}(\sigma_{x}-1/2)\sum_{\rho}\frac{1}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})}.$
 Using this bound and (3.6), we obtain
-$\sum_{\rho}\frac{1}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})}\leq\frac{4}{\sigma_{x}-1/2}\left|\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\right|+O\left(\frac{\log(|t|+M_{F})}{\sigma_{x}-1/2}\right).$ (3.10)
-On the other hand, by Lemma 3.1 and the bound (1.6), we have
-$\begin{split}\frac{L^{\prime}}{L}(\sigma+it,F)=&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}+\frac{w_{F}(x,\sigma,t)}{\log^{2}x}\sum_{\rho}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big{(}(\sigma-\beta)^{2}+(t-\gamma)^{2}\big{)}^{3/2}}+O\left(\frac{x^{5/14-\sigma}}{\log^{2}x}\right),\end{split}$ (3.11)
+$\sum_{\rho}\frac{1}{((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2})}\leq\frac{4}{\sigma_{x}-1/2}\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|+O\bigg(\frac{\log(|t|+M_{F})}{\sigma_{x}-1/2}\bigg).$ (3.10)
+On the other hand, by Lemma 3.1 and the estimate (1.6), we have, for $\sigma\geq\sigma_{x}$,
+$\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)=$ $-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}+\frac{w_{F}(x,\sigma,t)}{\log^{2}x}\sum_{\rho}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)^{3/2}}$
+$+O\bigg(\frac{x^{5/14-\sigma}}{\log^{2}x}\bigg),$ (3.11)
 with $|w_{F}(x,\sigma,t)|\leq 1$. Next we claim that
@@
 Using (3.10) and (3.12), we have
-[…]
-Inserting this bound into (3.11), we have
-$\begin{split}\frac{L^{\prime}}{L}(\sigma+it,F)\leq&-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}+\frac{4}{5}w^{\prime}_{F}(x,\sigma,t)x^{1/4-\sigma/2}\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\\ &+O\left(x^{1/4-\sigma/2}\log(|t|+M_{F})\right),\end{split}$ (3.13)
+$\begin{split}&\sum_{\rho}\frac{x^{\beta-\sigma}(1+x^{\beta-\sigma})^{2}}{\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)^{3/2}}\\ \leq&\,2\log x\sum_{\rho}\frac{x^{1/4-\sigma/2}}{(\beta-\sigma_{x})^{2}+(t-\gamma)^{2}}\\ \leq&\,8\log x\frac{x^{1/4-\sigma/2}}{\sigma_{x}-1/2}\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|+O\bigg(\frac{(\log x)x^{1/4-\sigma/2}\log(|t|+M_{F})}{\sigma_{x}-1/2}\bigg)\\ \leq&\,\frac{4}{5}(\log x)^{2}x^{1/4-\sigma/2}\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|+O\bigg((\log x)^{2}x^{1/4-\sigma/2}\log\big(|t|+M_{F}\big)\bigg).\end{split}$
+Inserting this bound into (3), we have
+$\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)=$ $-\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}+\frac{4}{5}w^{\prime}_{F}(x,\sigma,t)x^{1/4-\sigma/2}\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)$
+$+O\bigg(x^{1/4-\sigma/2}\log(|t|+M_{F})\bigg),$ (3.13)
 with $|w^{\prime}_{F}(x,\sigma,t)|\leq 1$. By taking $\sigma=\sigma_{x}$,
-$\begin{split}&\left(1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\right)\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\\ =&\,O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\left(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\right).\end{split}$
-Since $\left|1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\right|\geq 1-\left|\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\right|\geq\frac{1}{5}$, we have
-$\frac{L^{\prime}}{L}(\sigma_{x}+it,F)=O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\left(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\right).$ (3.14)
-Inserting (3.14) into (3.13) and (3.10), respectively, we complete the proof of lemma. ∎
+$\begin{split}&\bigg(1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\bigg)\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\\ =&\,O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\bigg(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\bigg).\end{split}$
+Since $\Big|1-\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\Big|\geq 1-\Big|\frac{4}{5}w^{\prime}_{F}(x,\sigma_{x},t)x^{1/4-\sigma_{x}/2}\Big|\geq\frac{1}{5}$, we have
+$\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)=O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\bigg(x^{1/4-\sigma_{x}/2}\log(|t|+M_{F})\bigg).$ (3.14)
+Then the claimed results follow from (3) and (3.14), and from (3.10) and (3.14), respectively. ∎
 The following theorem is to provide an approximation of $S_{F}(t)$.
 Theorem 3.4.
-For $t\neq 0$, and $x\geq 4$, we have
-$\begin{split}S_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}\log n}+O\left((\sigma_{x}-1/2)\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)\\ &+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)},\end{split}$
+For $t\neq 0$ and $x\geq 4$, we have
+$\begin{split}S_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}\log n}+O\bigg((\sigma_{x}-1/2)\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)\\ &+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big),\end{split}$
 where $\sigma_{x}$ is defined in (3.5).
@@
 By the definition of $S_{F}(t)$, we have
-$\begin{split}\pi S_{F}(t)=&-\int_{1/2}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+it,F)\mathrm{d}\sigma\\ =&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+it,F)\mathrm{d}\sigma-(\sigma_{x}-1/2)\operatorname{Im}\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\\ &+\int_{1/2}^{\sigma_{x}}\operatorname{Im}\left(\frac{L^{\prime}}{L}(\sigma_{x}+it,F)-\frac{L^{\prime}}{L}(\sigma+it,F)\right)\mathrm{d}\sigma\\ :=&\,J_{1}+J_{2}+J_{3}.\end{split}$
+$\begin{split}\pi S_{F}(t)=&-\int_{1/2}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\mathrm{d}\sigma\\ =&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\mathrm{d}\sigma-(\sigma_{x}-1/2)\operatorname{Im}\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\\ &+\int_{1/2}^{\sigma_{x}}\operatorname{Im}\bigg(\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)-\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\bigg)\mathrm{d}\sigma\\ :=&\,J_{1}+J_{2}+J_{3}.\end{split}$
 By Lemma 3.3, we have
-$\begin{split}J_{1}=&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+it,F)\mathrm{d}\sigma\\ =&\int_{\sigma_{x}}^{\infty}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}\mathrm{d}\sigma+O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\right)\\ &+O\left(\log(|t|+M_{F})\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\right)\\ =&\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}\log n}+O\left(\frac{1}{\log x}\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\left(\frac{\log(|t|+M_{F})}{\log x}\right).\end{split}$
+$\begin{split}J_{1}=&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\mathrm{d}\sigma\\ =&\int_{\sigma_{x}}^{\infty}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}\mathrm{d}\sigma+O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\bigg)\\ &+O\bigg(\log(|t|+M_{F})\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\bigg)\\ =&\,\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}\log n}+O\bigg(\frac{1}{\log x}\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\bigg(\frac{\log(|t|+M_{F})}{\log x}\bigg).\end{split}$
 Taking $\sigma=\sigma_{x}$ in Lemma 3.3, we have
-$\begin{split}J_{2}\leq&\,(\sigma_{x}-1/2)\left|\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\right|\\ \ll&\,(\sigma_{x}-1/2)\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|+(\sigma_{x}-1/2)\log(|t|+M_{F}).\end{split}$
+$\begin{split}J_{2}\leq&\,(\sigma_{x}-1/2)\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|\\ \ll&\,(\sigma_{x}-1/2)\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|+(\sigma_{x}-1/2)\log(|t|+M_{F}).\end{split}$
 Using Lemma 3.2, we get
-$\begin{split}&\operatorname{Im}\left(\frac{L^{\prime}}{L}(\sigma_{x}+it,F)-\frac{L^{\prime}}{L}(\sigma+it,F)\right)\\ =&\sum_{\rho}\frac{(\sigma_{x}-\sigma)(\sigma+\sigma_{x}-2\beta)(t-\gamma)}{\big{(}(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big{)}\big{(}(\sigma-\beta)^{2}+(t-\gamma)^{2}\big{)}}+O(1).\end{split}$
+$\begin{split}&\operatorname{Im}\bigg(\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)-\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\bigg)\\ =&\sum_{\rho}\frac{(\sigma_{x}-\sigma)(\sigma+\sigma_{x}-2\beta)(t-\gamma)}{\big((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big)\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)}+O(1).\end{split}$
 Hence
@@
 Thus,
-$\begin{split}\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma&\leq 3(\sigma_{x}-1/2)\int_{-\infty}^{\infty}\frac{|t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma\\ &\leq 10(\sigma_{x}-1/2).\end{split}$ (3.15)
+$\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma$ $\leq 3(\sigma_{x}-1/2)\int_{-\infty}^{\infty}\frac{|t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma$
+$\leq 10(\sigma_{x}-1/2).$ (3.15)
 If $|\beta-1/2|>\frac{1}{2}(\sigma_{x}-1/2)$, then
@@
 Thus,
-$\begin{split}\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma&<\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta|}{|t-\gamma|}\mathrm{d}\sigma\\ &\leq\int_{1/2}^{\sigma_{x}}\frac{6|\beta-1/2|}{3|\beta-1/2|}\mathrm{d}\sigma=2(\sigma_{x}-1/2).\end{split}$ (3.16)
-By the bounds (3.15)–(3.16) and Lemma 3.3, we obtain
-$\begin{split}|J_{3}|\leq&10(\sigma_{x}-1/2)\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\sigma_{x}-1/2)\\ =&O\left((\sigma_{x}-1/2)\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}.\end{split}$
-Thus we complete the proof of lemma. ∎
+$\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma$ $<\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta|}{|t-\gamma|}\mathrm{d}\sigma$
+$\leq\int_{1/2}^{\sigma_{x}}\frac{6|\beta-1/2|}{3|\beta-1/2|}\mathrm{d}\sigma=2(\sigma_{x}-1/2).$ (3.16)
+By the estimates in (3)–(3) and Lemma 3.3, we obtain
+$\begin{split}|J_{3}|\leq&10(\sigma_{x}-1/2)\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\sigma_{x}-1/2)\\ =&O\bigg((\sigma_{x}-1/2)\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big).\end{split}$
+Thus we complete the proof of the theorem. ∎
 4. Proof of Theorem 1.1
 Lemma 4.1.
-Let $t>0$ be fixed and $T$ be any large parameter. For $x=T^{\delta/3}$ with $0<\delta<\frac{3\theta}{8nk+3}$, $k\in\mathbb{N}$, we have
+Let $t>0$ be fixed and $T$ be any large positive parameter. For $k\in\mathbb{N}$, $x=T^{\delta/3}$ with $0<\delta<\frac{3\theta}{8nk+3}$, we have
 $\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{4n}x^{4nk(\sigma_{x}-1/2)}\ll_{t,n,\delta,k}\frac{1}{(\log T)^{4n}},\end{split}$
@@
 For $k\in\mathbb{N}$, by the definition of $\sigma_{x,F}$, we have
-$\begin{split}&(\sigma_{x,F}-1/2)^{4n}x^{4nk(\sigma_{x,F}-1/2)}\\ \leq&\left(\frac{10}{\log x}\right)^{4n}x^{40nk/\log x}+2^{4n+1}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}.\end{split}$
+$\begin{split}&(\sigma_{x,F}-1/2)^{4n}x^{4nk(\sigma_{x,F}-1/2)}\\ \leq&\bigg(\frac{10}{\log x}\bigg)^{4n}x^{40nk/\log x}+2^{4n+1}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}.\end{split}$
 On the other hand,
-$\begin{split}&\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\left(\frac{i+1}{\log x}\right)^{4n}x^{8nk\frac{i+1}{\log x}}\sum_{\frac{1}{2}+\frac{i}{\log x}<\beta\leq\frac{1}{2}+\frac{i+1}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}1\\ \leq&\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\left(i+1\right)^{4n}e^{8nk(i+1)}N\left(\frac{i}{\log x},|t|+\frac{e^{3(i+1)}}{\log x};F\right).\end{split}$
+$\begin{split}&\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\bigg(\frac{i+1}{\log x}\bigg)^{4n}x^{8nk\frac{i+1}{\log x}}\sum_{\frac{1}{2}+\frac{i}{\log x}<\beta\leq\frac{1}{2}+\frac{i+1}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}1\\ \leq&\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\big(i+1\big)^{4n}\mathrm{e}^{8nk(i+1)}N\bigg(\frac{i}{\log x},|t|+\frac{\mathrm{e}^{3(i+1)}}{\log x};F\bigg).\end{split}$
 By Lemma 2.4, we have
-$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\left(i+1\right)^{4n}e^{8nk(i+1)}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}N\left(\frac{i}{\log x},|t|+\frac{e^{3(i+1)}}{\log x};F\right)\\ \ll&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\infty}\left(i+1\right)^{4n}e^{8nk(i+1)}\left(|t|+\frac{e^{3(i+1)}}{\log x}\right)T^{-\theta\frac{i}{\log x}}\log T\\ \ll&{}_{t,n,\delta,k}\,\frac{\log T}{(\log x)^{4n+1}}\sum_{i=5}^{\infty}\left(i+1\right)^{4n}e^{(8nk+3-\frac{3\theta}{\delta})i}\\ \ll&{}_{t,n,\delta,k}\,\frac{1}{(\log T)^{4n}}\end{split}$
-provided that $0<\delta<\frac{3\theta}{8nk+3}$. In addition, by Corollary 2.3,
-$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\frac{10}{\log x}\right)^{4n}x^{40nk/\log x}\ll_{n,k}\frac{1}{(\log x)^{4n}}\ll_{n,k,\delta}\frac{1}{(\log T)^{4n}}.\end{split}$
-Thus we complete the proof of lemma. ∎
+$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\big(i+1\big)^{4n}\mathrm{e}^{8nk(i+1)}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}N\bigg(\frac{i}{\log x},|t|+\frac{\mathrm{e}^{3(i+1)}}{\log x};F\bigg)\\ \ll&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\infty}\big(i+1\big)^{4n}\mathrm{e}^{8nk(i+1)}\bigg(|t|+\frac{\mathrm{e}^{3(i+1)}}{\log x}\bigg)T^{-\theta\frac{i}{\log x}}\log T\\ \ll&{}_{t,n,\delta,k}\,\frac{\log T}{(\log x)^{4n+1}}\sum_{i=5}^{\infty}\big(i+1\big)^{4n}\mathrm{e}^{(8nk+3-\frac{3\theta}{\delta})i}\\ \ll&{}_{t,n,\delta,k}\,\frac{1}{(\log T)^{4n}}\end{split}$
+provided that $0<\delta<\frac{3\theta}{8nk+3}$. In addition, by Corollary 2.3 and (1.9),
+$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\frac{10}{\log x}\bigg)^{4n}x^{40nk/\log x}\ll_{n,k}\frac{1}{(\log x)^{4n}}\ll_{n,k,\delta}\frac{1}{(\log T)^{4n}}.\end{split}$
+Thus we complete the proof of the lemma. ∎
 For a positive parameter $x$ (to be chosen later), let
-$\mathcal{M}_{F}(t):=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+it}}\qquad\text{and}\qquad\mathcal{R}_{F}(t):=S_{F}(t)-\mathcal{M}_{F}(t),$ (4.1)
+$\mathcal{M}_{F}(t):=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+\mathrm{i}t}}\qquad\text{and}\qquad\mathcal{R}_{F}(t):=S_{F}(t)-\mathcal{M}_{F}(t),$ (4.1)
 where we recall (3.2) that $C_{F}(p)=\alpha_{F,1}(p)+\alpha_{F,2}(p)+\alpha_{F,3}(p)=A_{F}(1,p)$.
@@
 We have
-$\begin{split}\mathcal{R}_{F}(t)=&O\left(\left|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}\right|\right)+O\left(\sum_{k=2}^{7}\left|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}\right|\right)\\ &+O\left((\sigma_{x}-1/2)\sum_{k=1}^{7}x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)\\ &+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}+O\big{(}1\big{)}.\end{split}$
+$\begin{split}\mathcal{R}_{F}(t)=&O\bigg(\bigg|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}\bigg|\bigg)+O\bigg(\sum_{k=2}^{7}\bigg|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}\bigg|\bigg)\\ &+O\bigg((\sigma_{x}-1/2)\sum_{k=1}^{7}x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)\\ &+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big)+O\big(1\big).\end{split}$
 Proof.
 By Theorem 3.4,
-$\begin{split}\mathcal{R}_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)(\Lambda_{x}(p)p^{1/2-\sigma_{x}}-\Lambda(p))}{p^{1/2+it}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}\log p^{k}}\\ &+O\left((\sigma_{x}-1/2)\left|\sum_{k=1}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}\right|\right)+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}.\end{split}$
+$\begin{split}\mathcal{R}_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)(\Lambda_{x}(p)p^{1/2-\sigma_{x}}-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}\log p^{k}}\\ &+O\bigg((\sigma_{x}-1/2)\bigg|\sum_{k=1}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}\bigg|\bigg)+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big).\end{split}$
 Using the bound (1.6), we deduce that
-$\begin{split}\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}=O(1),\quad\text{and}\quad\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}\log p}=O(1).\end{split}$
+$\begin{split}\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}=O(1),\quad\text{and}\quad\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}\log p}=O(1).\end{split}$
 Thus,
-$\begin{split}\mathcal{R}_{F}(t)=&\,\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p^{k}}\\ &+\frac{1}{\pi}\operatorname{Im}\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+it)}\log p^{k}}\\ &+O\left((\sigma_{x}-1/2)\left|\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}\right|\right)\\ &+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}+O(1).\end{split}$
+$\begin{split}\mathcal{R}_{F}(t)=&\,\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p^{k}}\\ &+\frac{1}{\pi}\operatorname{Im}\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+\mathrm{i}t)}\log p^{k}}\\ &+O\bigg((\sigma_{x}-1/2)\bigg|\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}\bigg|\bigg)\\ &+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big)+O(1).\end{split}$
 Note that for $1/2\leq a\leq\sigma_{x}$ and $k\geq 1$,
-$\begin{split}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+it)}}\right|&=kx^{k(a-1/2)}\left|\int_{a}^{\infty}x^{k(1/2-\sigma)}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\mathrm{d}\sigma\right|\\ &\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma.\end{split}$
+$\begin{split}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+\mathrm{i}t)}}\bigg|&=kx^{k(a-1/2)}\bigg|\int_{a}^{\infty}x^{k(1/2-\sigma)}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\mathrm{d}\sigma\bigg|\\ &\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma.\end{split}$
 Thus,
-$\begin{split}&\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+it)}\log p^{k}}\right|=\left|\int_{1/2}^{\sigma_{x}}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+it)}}\mathrm{d}a\right|\\ &\leq k(\sigma_{x}-1/2)x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma,\end{split}$
+$\begin{split}&\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+\mathrm{i}t)}\log p^{k}}\bigg|=\bigg|\int_{1/2}^{\sigma_{x}}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+\mathrm{i}t)}}\mathrm{d}a\bigg|\\ &\leq k(\sigma_{x}-1/2)x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma,\end{split}$
 and
-$\begin{split}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}\right|\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma.\end{split}$
-Thus we complete the proof of lemma. ∎
+$\begin{split}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}\bigg|\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma.\end{split}$
+Thus we complete the proof of the lemma. ∎
 Lemma 4.3.
-Let $t>0$ be fixed and $T$ be any large parameter. For $x=T^{\delta/3}$ with sufficiently small $\delta>0$, we have
+Let $t>0$ be fixed and $T$ be any large positive parameter. For $x=T^{\delta/3}$ with sufficiently small $\delta>0$, we have
 $\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\mathcal{M}_{F}(t)^{n}=C_{n}(\log\log T)^{\frac{n}{2}}+O_{t,n}\big((\log\log T)^{\frac{n-1}{2}}\big)\end{split}$ (4.2)
@@
 Recall that
-$\mathcal{M}_{F}(t)=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+it}}=\frac{-i}{2\pi}\left(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)}{p^{1/2+it}}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-it}}\right).$
+$\mathcal{M}_{F}(t)=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+\mathrm{i}t}}=\frac{-\mathrm{i}}{2\pi}\bigg(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)}{p^{1/2+\mathrm{i}t}}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-\mathrm{i}t}}\bigg).$
 Set $x=T^{\delta/3}$ for a suitably small $\delta>0$. Thus we have that the general term of
-$\mathcal{M}_{F}(t)^{n}=\frac{(-i)^{n}}{(2\pi)^{n}}\left(\sum_{p\leq T^{\delta}}\frac{A_{F}(1,p)}{p^{1/2+it}}-\sum_{p\leq T^{\delta}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-it}}\right)^{n}$
+$\mathcal{M}_{F}(t)^{n}=\frac{(-\mathrm{i})^{n}}{(2\pi)^{n}}\bigg(\sum_{p\leq T^{\delta}}\frac{A_{F}(1,p)}{p^{1/2+\mathrm{i}t}}-\sum_{p\leq T^{\delta}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-\mathrm{i}t}}\bigg)^{n}$
 has the form
-$\frac{A_{F}(1,p_{1})^{m(p_{1})}\overline{A_{F}(1,p_{1})^{n(p_{1})}}}{p_{1}^{m(p_{1})(1/2+it)}(-1)^{n(p_{1})}p_{1}^{n(p_{1})(1/2-it)}}\times\cdots\times\frac{A_{F}(1,p_{r})^{m(p_{r})}\overline{A_{F}(1,p_{r})^{n(p_{r})}}}{p_{r}^{m(p_{r})(1/2+it)}(-1)^{n(p_{r})}p_{r}^{n(p_{r})(1/2-it)}},$ (4.4)
+$\frac{A_{F}(1,p_{1})^{m(p_{1})}\overline{A_{F}(1,p_{1})^{n(p_{1})}}}{p_{1}^{m(p_{1})(1/2+\mathrm{i}t)}(-1)^{n(p_{1})}p_{1}^{n(p_{1})(1/2-\mathrm{i}t)}}\times\cdots\times\frac{A_{F}(1,p_{r})^{m(p_{r})}\overline{A_{F}(1,p_{r})^{n(p_{r})}}}{p_{r}^{m(p_{r})(1/2+\mathrm{i}t)}(-1)^{n(p_{r})}p_{r}^{n(p_{r})(1/2-\mathrm{i}t)}},$ (4.4)
 where $p_{1}<p_{2}<\cdots<p_{r}\leq T^{\delta}$, $m(p_{j})+n(p_{j})\geq 1$, and $\sum_{j=1}^{r}(m(p_{j})+n(p_{j}))=n$.
 Now we discuss the contribution from the general term (4.4) in the following three cases.
-[…]
+Case (I): In the general term (4.4), $m(p_{j_{0}})\not\equiv n(p_{j_{0}})(\bmod\,3)$ for some $j_{0}$. By Lemma 2.1, we have
 $\begin{split}A_{F}(1,p_{j_{0}})^{m(p_{j_{0}})}=\sum_{k,\ell\geq 0\atop k+\ell\leq m(p_{j_{0}})}a_{k,\ell}(m(p_{j_{0}}))A_{F}(p_{j_{0}}^{k},p_{j_{0}}^{\ell})\end{split}$
@@
 There equations would further lead to
-[…]
+$\begin{split}m(p_{j_{0}})-n(p_{j_{0}})&=(m_{1}-n_{1})+(m_{2}-n_{2})+(m_{3}-n_{3})\\ &=3(m_{2}-n_{2})\equiv 0(\bmod\,3),\end{split}$
 which contradicts the assumption for Case (I).
 With the above observation and the Hecke relations, the numerator of this general term (4.4) can be written as a sum whose terms are all of the form $A_{F}(m_{1},n_{1})\overline{A_{F}(m_{2},n_{2})}$ where $(m_{1},n_{1})\neq(m_{2},n_{2})$.
-Since for these terms $\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}=0$, we see from Theorem 2.2 and the estimate (1.8) that the contribution from the general term (4.4) to the moment (4.2) is
-$\begin{split}&\ll_{n}\frac{1}{\mathcal{H}}\sum_{p_{1}<p_{2}<\cdots<p_{r}\leq T^{\delta}}p_{1}^{-\frac{1}{2}(m(p_{1})+n(p_{1}))}\cdots p_{r}^{-\frac{1}{2}(m(p_{r})+n(p_{r}))}\\ &\qquad\qquad\qquad\qquad\qquad\times\left((T^{1+2n\delta})^{\varepsilon}\big{(}T^{1+n\delta}+T^{3}+T^{1+2n\delta\vartheta}M^{2}\big{)}\right)\\ &\ll_{n}\frac{1}{\mathcal{H}}\bigg{(}\sum_{p\leq T^{\delta}}p^{-\frac{1}{2}}\bigg{)}^{r}\big{(}T^{1+\varepsilon+n\delta(1+2\varepsilon)}+T^{3+\varepsilon+2n\delta\varepsilon}+T^{1+\varepsilon+2n\delta(\vartheta+\varepsilon)}M^{2}\big{)}\\ &\ll_{n}T^{-2+\varepsilon+n\delta(3/2+2\varepsilon)}M^{-2}+T^{\varepsilon+n\delta(1/2+2\varepsilon)}M^{-2}+T^{-2+\varepsilon+n\delta(1/2+2\vartheta+2\varepsilon)},\end{split}$
+Since for these terms $\delta_{m_{1},m_{2}}\delta_{n_{1},n_{2}}=0$, we see from Theorem 2.2 and the estimate (1.9) that the contribution from the general term (4.4) to the moment (4.2) is
+$\begin{split}&\ll_{n}\frac{1}{\mathcal{H}}\sum_{p_{1}<p_{2}<\cdots<p_{r}\leq T^{\delta}}p_{1}^{-\frac{1}{2}(m(p_{1})+n(p_{1}))}\cdots p_{r}^{-\frac{1}{2}(m(p_{r})+n(p_{r}))}\\ &\qquad\qquad\qquad\qquad\qquad\times\bigg((T^{1+2n\delta})^{\varepsilon}\big(T^{1+n\delta}+T^{3}+T^{1+2n\delta\vartheta}M^{2}\big)\bigg)\\ &\ll_{n}\frac{1}{\mathcal{H}}\bigg(\sum_{p\leq T^{\delta}}p^{-\frac{1}{2}}\bigg)^{r}\big(T^{1+\varepsilon+n\delta(1+2\varepsilon)}+T^{3+\varepsilon+2n\delta\varepsilon}+T^{1+\varepsilon+2n\delta(\vartheta+\varepsilon)}M^{2}\big)\\ &\ll_{n}T^{-2+\varepsilon+n\delta(3/2+2\varepsilon)}M^{-2}+T^{\varepsilon+n\delta(1/2+2\varepsilon)}M^{-2}+T^{-2+\varepsilon+n\delta(1/2+2\vartheta+2\varepsilon)},\end{split}$
 which is negligible upon taking $0<\delta<\frac{1}{n}$.
-[…]
+Case (II): In the general term (4.4), $m(p_{j})\equiv n(p_{j})(\bmod\,3)$ for all $j$, and $m(p_{j_{0}})\geq 2$ or $n(p_{j_{0}})\geq 2$ for some $j_{0}$.
+In this case, $m(p_{j})+n(p_{j})\geq 2$ for all $j$ and $m(p_{j_{0}})+n(p_{j_{0}})\geq 3$ for some $j_{0}$, so $n>2r$ and $r<\lfloor\frac{n-1}{2}\rfloor$.
+By applying Lemma 2.1 to the numerator of this general term (4.4) and multiplying out by Hecke relations, one sees that it is possible to have resulting terms $A_{F}(m_{1},n_{1})\overline{A_{F}(m_{2},n_{2})}$ where $(m_{1},n_{1})=(m_{2},n_{2})$.
+Using Theorem 2.2, the contribution to the moment (4.2) such terms is bounded by the diagonal term of size $\asymp T^{3}M^{2}$.
+The contribution from the off-diagonal terms is clearly negligible compared to $T^{3}M^{2}$, given that $0<\delta<\frac{1}{n}$.
+Using the classical result
 $\sum_{p\leq x}\frac{1}{p}=\log\log x+O(1),$ (4.5)
@@
 Case (III): $m(p_{j})=n(p_{j})=1$ for all $j$. The contribution from these terms to (4.2) is
-$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\binom{2m}{m}(m!)(m!)\frac{(-1)^{m}(-i)^{2m}}{(2\pi)^{2m}}\\ &\qquad\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{A_{F}(1,p_{1})\overline{A_{F}(1,p_{1})}\cdots A_{F}(1,p_{m})\overline{A_{F}(1,p_{m})}}{p_{1}\cdots p_{m}}.\end{split}$ (4.6)
+$\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\binom{2m}{m}(m!)(m!)\frac{(-1)^{m}(-\mathrm{i})^{2m}}{(2\pi)^{2m}}$
+$\qquad\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{A_{F}(1,p_{1})\overline{A_{F}(1,p_{1})}\cdots A_{F}(1,p_{m})\overline{A_{F}(1,p_{m})}}{p_{1}\cdots p_{m}}.$ (4.6)
 By the Hecke relations, we have
 $\begin{split}&A_{F}(1,p_{1})\overline{A_{F}(1,p_{1})}\cdots A_{F}(1,p_{m})\overline{A_{F}(1,p_{m})}\\ &\quad=A_{F}(1,p_{1}\cdots p_{m})\overline{A_{F}(1,p_{1}\cdots p_{m})}.\end{split}$
-Applying Theorem 2.2 and the estimate (4.5), we have (4.6) equals
-$\begin{split}&\frac{1}{\mathcal{H}}\frac{(2m)!}{(2\pi)^{2m}}\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{1}{p_{1}\cdots p_{m}}\cdot\left(\mathcal{H}+O\left(T^{1+\varepsilon}(p_{1}\cdots p_{m})+T^{3+\varepsilon}+T^{1+\varepsilon}M^{2}(p_{1}\cdots p_{m})^{2\vartheta}\right)\right)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}+O\bigg{(}T^{-2+\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}1\bigg{)}\\ &+O\bigg{(}T^{\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}\bigg{)}+O\bigg{(}T^{-2+\varepsilon}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{(p_{1}\cdots p_{m})^{1-2\vartheta}}\bigg{)}\\ \\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T+O(1))^{m}+O\big{(}T^{m\delta-2+\varepsilon}M^{-2}+T^{\varepsilon}M^{-2}(\log\log T)^{m}+T^{2m\delta\vartheta-2+\varepsilon}(\log\log T)^{m}\big{)}\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T)^{m}+O_{m}\big{(}(\log\log T)^{m-1}\big{)}.\end{split}$
+Applying Theorem 2.2 and the estimate (4.5), we have (4) equals
+$\begin{split}&\frac{1}{\mathcal{H}}\frac{(2m)!}{(2\pi)^{2m}}\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{1}{p_{1}\cdots p_{m}}\cdot\bigg(\mathcal{H}+O\bigg(T^{1+\varepsilon}(p_{1}\cdots p_{m})+T^{3+\varepsilon}+T^{1+\varepsilon}M^{2}(p_{1}\cdots p_{m})^{2\vartheta}\bigg)\bigg)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}+O\bigg(T^{-2+\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}1\bigg)\\ &+O\bigg(T^{\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}\bigg)+O\bigg(T^{-2+\varepsilon}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{(p_{1}\cdots p_{m})^{1-2\vartheta}}\bigg)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T+O(1))^{m}+O\big(T^{m\delta-2+\varepsilon}M^{-2}+T^{\varepsilon}M^{-2}(\log\log T)^{m}+T^{2m\delta\vartheta-2+\varepsilon}(\log\log T)^{m}\big)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T)^{m}+O_{m}\big((\log\log T)^{m-1}\big).\end{split}$
 Now the asymptotic formula (4.2) follows from Cases (I)–(III).
 Next we are ready to prove (4.3). Recall that $\sigma_{x}=\sigma_{x,F}$ depending on $F$. By Lemma 4.2,
-$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}\\ \ll&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}\right|^{2n}+\sum_{k=2}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}\right|^{2n}\\ &+\sum_{k=1}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{2n}\\ &+\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}\big{(}\log(|t|+M_{F})\big{)}^{2n}+\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}.\end{split}$ (4.7)
-For the first two terms in (4.7), we have
-$\begin{split}&\left|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}\right|^{2n}\ll_{n}\left(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2-it}\log p}\right)^{2n},\end{split}$
+$\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}$
+$\ll$ $\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}\bigg|^{2n}+\sum_{k=2}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}\bigg|^{2n}$
+$+$ $\sum_{k=1}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{2n}$
+$+$ $\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}\big(\log(|t|+M_{F})\big)^{2n}+\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}.$ (4.7)
+For the first two terms in (4), we have
+$\begin{split}&\bigg|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}\bigg|^{2n}\\ &\qquad\ll_{n}\bigg(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2-\mathrm{i}t}\log p}\bigg)^{2n},\end{split}$
 and for $k\geq 2$,
-$\begin{split}\left|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}\right|^{2n}\ll_{n}\left(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})}{p^{k(1/2-it)}\log p}\right)^{2n}.\end{split}$ (4.8)
+$\begin{split}\bigg|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}\bigg|^{2n}\ll_{n}\bigg(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})}{p^{k(1/2-\mathrm{i}t)}\log p}\bigg)^{2n}.\end{split}$ (4.8)
 For (4.8), we note that
 $\begin{split}C_{F}(p^{k})=\alpha_{F,1}(p)^{k}+\alpha_{F,2}(p)^{k}+\alpha_{F,3}(p)^{k}\end{split}$
-is a symmetric polynomial in $\alpha_{F,j}(p)$, since (1.5) says that $1$, $A_{F}(1,p)$, $A_{F}(p,1)=\overline{A_{F}(1,p)}$ are the elementary symmetric polynomials in $\alpha_{F,j}(p)$, the fundamental theorem of symmetric polynomials (see [15, pg.
+is a symmetric polynomial in $\alpha_{F,j}(p)$, since (1.5) says that $1$, $A_{F}(1,p)$, $A_{F}(p,1)=\overline{A_{F}(1,p)}$ are the elementary symmetric polynomials in $\alpha_{F,j}(p)$, the fundamental theorem of symmetric polynomials (see [17, pg.
 21, Remark]) implies that the symmetric polynomial $C_{F}(p^{k})=\alpha_{F,1}(p)^{k}+\alpha_{F,2}(p)^{k}+\alpha_{F,3}(p)^{k}$ is a polynomial in $1$, $A_{F}(1,p)$ and $\overline{A_{F}(1,p)}$.
-Then we can use the fact $|\Lambda_{x}(p)-\Lambda(p)|=O\left(\frac{\log^{3}p}{\log^{2}x}\right)$ and the similar argument as Case (I)–(III) to see that the contributions from first two terms in (4.7) to (4.3) is of $O_{t,n}(1)$.
-The contributions from last two terms in (4.7) is of $O_{t,n}(1)$ by Cauchy–Schwarz inequality, the assumption (1.7), Corollary 2.3 and Lemma 4.1.
-For the third term in (4.7), it follows from Cauchy–Schwarz inequality that
-$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{2n}\\ &\leq\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{4n}x^{4nk(\sigma_{x}-1/2)}\right)^{1/2}\\ &\times\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{4n}\right)^{1/2},\end{split}$ (4.9)
-where $1\leq k\leq 7$. By Hölder’s inequality with the exponents $4n/(4n-1)$ and $4n$,
-$\begin{split}&\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{4n}\\ \leq&\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\mathrm{d}\sigma\right)^{4n-1}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}\mathrm{d}\sigma\right)\\ =&\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}\mathrm{d}\sigma.\end{split}$
+Then we can use the fact $|\Lambda_{x}(p)-\Lambda(p)|=O\big(\frac{\log^{3}p}{\log^{2}x}\big)$ and the similar argument as Case (I)–(III) to see that the contributions from first two terms in (4) to (4.3) are of $O_{t,n}(1)$.
+The contributions from last two terms in (4) are of $O_{t,n}(1)$ by Cauchy–Schwarz inequality, the assumption (1.8), Corollary 2.3 and Lemma 4.1.
+For the third term in (4), it follows from Cauchy–Schwarz inequality that
+$\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{2n}$
+$\leq\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{4n}x^{4nk(\sigma_{x}-1/2)}\bigg)^{1/2}$
+$\times\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{4n}\bigg)^{1/2},$ (4.9)
+where $1\leq k\leq 7$. By Hölder’s inequality with the exponents 4n/(4n1)4n/(4n-1) and $4n$,
+$\begin{split}&\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{4n}\\ \leq&\,(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\mathrm{d}\sigma\bigg)^{4n-1}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}\mathrm{d}\sigma\bigg)\\ =&\,\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}\mathrm{d}\sigma.\end{split}$
 Note that
-$\begin{split}&\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}\\ &\ll_{n}\left(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}+\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-it)}}\right)^{4n}\\ &\quad+\left(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-it)}}\right)^{4n}.\end{split}$
+$\begin{split}&\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}\\ &\ll_{n}\bigg(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}+\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-\mathrm{i}t)}}\bigg)^{4n}\\ &\quad+\bigg(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-\mathrm{i}t)}}\bigg)^{4n}.\end{split}$
 By using the same argument as in proving (4.2), for $1\leq k\leq 7$, we get
-$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}&\ll_{t,n}(\log x)^{8n}.\end{split}$
+$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}&\ll_{t,n}(\log x)^{8n}.\end{split}$
 Thus,
-$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{4n}\\ \ll_{t,n}\,&\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}(\log x)^{8n}\mathrm{d}\sigma\ll_{t,n}(\log x)^{4n}.\end{split}$
-Inserting this bound to (4.9) and using Lemma 4.1, we have the third term in (4.7) is of $O_{t,n}(1)$. Thus we complete the proof of this lemma. ∎
+$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{4n}\\ &\ll_{t,n}\,\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}(\log x)^{8n}\mathrm{d}\sigma\ll_{t,n}(\log x)^{4n}.\end{split}$
+Inserting this bound to (4) and using Lemma 4.1, we have the third term in (4) is of $O_{t,n}(1)$. This completes the proof of the lemma. ∎
 Proof of Theorem 1.1. By the binomial theorem, we have
-$\begin{split}S_{F}(t)^{n}=\mathcal{M}_{F}(t)^{n}+O_{n}\left(\sum_{\ell=1}^{n}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\right),\end{split}$ (4.10)
+$\begin{split}S_{F}(t)^{n}=\mathcal{M}_{F}(t)^{n}+O_{n}\bigg(\sum_{\ell=1}^{n}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\bigg),\end{split}$ (4.10)
 where $\mathcal{M}_{F}(t)$ and $\mathcal{R}_{F}(t)$ are defined in (4.1). For $1\leq\ell\leq n$, we apply the generalized Hölder’s inequality, Corollary 2.3 and Lemma 4.3 to get
-$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\\ \ll&\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\right)^{\frac{1}{2}}\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{2n}\right)^{\frac{n-\ell}{2n}}\\ &\times\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}\right)^{\frac{\ell}{2n}}\\ \ll&{}_{t,n}\,(\log\log T)^{(n-\ell)/2}\ll_{t,n}(\log\log T)^{(n-1)/2}.\end{split}$
-Here we use the facts that the positivity of $\mathcal{N}_{F}$ and $h_{T,M}(\mu_{F})\geq 0$ for all $\rm GL_{3}$ Hecke–Maass cusp forms $F$.
+$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\\ \ll&\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg)^{\frac{1}{2}}\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{2n}\bigg)^{\frac{n-\ell}{2n}}\\ &\times\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}\bigg)^{\frac{\ell}{2n}}\\ \ll&{}_{t,n}\,(\log\log T)^{(n-\ell)/2}\ll_{t,n}(\log\log T)^{(n-1)/2}.\end{split}$
+Here we have used the positivity of $\mathcal{N}_{F}$ and the fact that $h_{T,M}(\mu_{F})\geq 0$ for all $\mathrm{GL}_{3}$ Hecke–Maass cusp forms $F$.
 Then (1.11) follows from (4.2), (4.10) and the above estimate.
 For the second assertion of Theorem 1.1, note that the $n$-th moment of $\mu_{TM}$ is
-$\int_{\mathbb{R}}\xi^{n}\mathrm{d}\mu_{TM}(\xi)=\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\frac{S_{F}(t)}{\sqrt{\log\log T}}\right)^{n}\right)\bigg{/}\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\right).$
+$\int_{\mathbb{R}}\xi^{n}\mathrm{d}\mu_{TM}(\xi)=\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\frac{S_{F}(t)}{\sqrt{\log\log T}}\bigg)^{n}\bigg)\bigg/\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg).$
 Using (1.11) and Corollary 2.3, we obtain that for all $n\geq 0$,
 $\lim_{T\rightarrow\infty}\int_{\mathbb{R}}\xi^{n}\mathrm{d}\mu_{TM}(\xi)=C_{n}=\sqrt{\pi}\int_{\mathbb{R}}\xi^{n}\exp(-\pi^{2}\xi^{2})\mathrm{d}\xi.$
-Then by the theory of moments in probability theory (see, for example, [2, Lemma 30.2]), we complete the proof of Theorem 1.1.
+Then by the theory of moments in probability theory (see, for example, [3, Theorem 30.2]), we complete the proof of Theorem 1.1.
+5. Conclusion
+In conclusion, we establish an unconditional asymptotic formula for the moments of $S_{F}(t)$, where $F$ is a Hecke–Maass cusp form for $\mathrm{SL}_{3}(\mathbb{Z})$ with Langlands parameter in generic position.
+Our main result, Theorem 1.1, extends the conditional result of Liu and Liu [15], which assumed the GRH, to an unconditional setting.
+This achievement is made possible by applying a novel weighted zero-density estimate for $L(s,F)$ in the spectral aspect, recently established in [20].
+Moreover, we demonstrate that the normalized values $S_{F}(t)/\sqrt{\log\log T}$ converge in distribution to a Gaussian random variable as $T\to\infty$.
+A central goal of our ongoing work is to extend this unconditional framework to $\mathrm{SL}_{n}(\mathbb{Z})$ for all $n\geq 3$, thereby providing an unconditional counterpart to the conditional theorem of Chen et al.
+[9].
+Acknowledgements Q.
+Sun was partially supported by the National Natural Science Foundation of China (Grant Nos.
+12471005 and 12031008) and the Natural Science Foundation of Shandong Province (Grant No.
+ZR2023MA003).
+The authors are very grateful to the referees for their valuable suggestions.